### v4

# Positivstellensätze for polynomial matrices with universal quantifiers

Abstract.
This paper investigates Positivstellensätze for polynomial matrices subject to universally quantified polynomial matrix inequality constraints. We first establish a matrix-valued Positivstellensatz under the Archimedean condition, incorporating universal quantifiers. For scalar-valued polynomial objectives, we further develop a sparse Positivstellensatz that leverages correlative sparsity patterns within these quantified constraints. Moving beyond the Archimedean framework, we then derive two generalized Positivstellensätze under analogous settings. These results collectively unify and extend foundational theorems in three distinct contexts: classical polynomial Positivstellensätze, their universally quantified counterparts, and matrix polynomial formulations. Applications of the established Positivstellensätze to robust polynomial matrix inequality constrained optimization are also investigated.
Key words and phrases:
Positivstellensatz, sum of squares, polynomial matrix inequality, polynomial matrix optimization, universal quantifier, correlative sparsity
1991 Mathematics Subject Classification:
90C23, 15A54, 13J30, 14P10, 11E25, 12D15
1. Introduction
Positivstellensätze are fundamental results in real algebraic geometry, asserting under which conditions a polynomial is guaranteed to be positive on a given set [5, 26, 31, 41]. Most Positivstellensätze achieve this by expressing the polynomial using sums of squares (SOS). These powerful results offer constructive methods for certifying whether a polynomial $f$ is positive over a basic semialgebraic set defined by polynomial inequalities
$\mathbf{K}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid g_{j}({\boldsymbol{x}})\geq 0,j=1,\ldots,s\},$
and have widespread applications in areas such as optimization, algebraic geometry, control theory, and more where polynomial positivity is a key concern [4, 14, 25, 26, 34].
Depending on whether the semialgebraic set $\mathbf{K}$ is assumed to be compact, Positivstellensätze are divided into two categories. When $\mathbf{K}$ is a compact polyhedron with non-empty interior, Handelman’s Positivstellensatz [12] states that a polynomial $f$ that is positive on $\mathbf{K}$ can be expressed as a positive linear combination of cross-products of $g_{j}$’s. In the case of a more general compact basic semialgebraic set $\mathbf{K}$, Schmüdgen [43] demonstrated that if a polynomial $f$ is positive on $\mathbf{K}$, then it belongs to the preorder generated by $g_{j}$’s, that is, $f$ can be represented as an SOS-weighted combination of cross-products of $g_{j}$’s. Putinar’s Positivstellensatz [39] provides an alternative representation that avoids the need for cross-products of $g_{j}$’s under the Archimedean condition (slightly stronger than compactness). Building on Putinar’s Positivstellensatz, Lasserre [22, 24, 25] introduced a moment-SOS hierarchy of semidefinite relaxations for polynomial optimization. This framework generates a non-decreasing sequence of lower bounds converging to the optimal value of a given polynomial optimization problem.
There have been several attempts to provide certificates of positivity of a polynomial over a non-compact semialgebraic set. Without requiring compactness of $\mathbf{K}$, Krivine-Stengle Positivstellensatz [21, 46] states that a polynomial $f$ is positive over $\mathbf{K}$ if and only if $\psi f=1+\phi$ for some $\psi,\phi$ from the preorder generated by $g_{j}$’s. Pólya [35] demonstrated that if $f$ is homogeneous and positive on ${\mathbb{R}}^{n}_{+}\setminus\{\mathbf{0}\}$, then multiplying $f$ by some power of $\sum_{i=1}^{n}x_{i}$, one obtains a polynomial with nonnegative coefficients. Dickinson and Povh [7] generalized Pólya’s Positivstellensatz for homogeneous polynomials being positive on the intersection ${\mathbb{R}}^{n}_{+}\cap\mathbf{K}\setminus\{\mathbf{0}\}$, where $g_{j}$’s are assumed to be also homogeneous. Reznick [40] proved that, after multiplying by some power of $\sum_{i=1}^{n}x_{i}^{2}$, any positive definite (PD) form is a sum of even powers of linear forms. Putinar and Vasilescu [38] extended Reznick’s result to the constrained case where $f$ and $g_{j}$’s are homogeneous polynomials of even degree. When $\mathbf{K}$ is non-compact, SOS-structured positivity certificates of $f$ over $\mathbf{K}$ could be also established by investigating various geometric objects associated with the data, e.g., gradient varieties [33], principal gradient tentacles [45], truncated tangency varieties [48], and the polar varieties [10].
Recently, Hu, Klep, and Nie [16] studied Positivstellensätze concerning semialgebraic sets defined by universal quantifiers (UQ). Specifically, for a given tuple $g=(g_{1},\ldots,g_{s})$ of real polynomials in $\boldsymbol{x}$} and ${\boldsymbol{y}}=(y_{1},\ldots,y_{m})$, and a closed set $\mathcal{Y}\subset{\mathbb{R}}^{m}$, they consider representations of polynomials that are positive over the semialgebraic set
$\mathcal{U}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid g_{1}({\boldsymbol{x}},{\boldsymbol{y}})\geq 0,\ldots,g_{s}({\boldsymbol{x}},{\boldsymbol{y}})\geq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\}.$
For a fixed measure $\nu$ with support exactly on $\mathcal{Y}$, under the Carleman condition on $\nu$ and the Archimedean condition, they proved that if a polynomial $f$ is positive on $\mathcal{U}$, then $f$ belongs to the quadratic module associated to $(g,\nu)$. In other words, $f$ admits a representation
$\sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\sigma_{1}({\boldsymbol{x}},{\boldsymbol{y}})g_{1}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})+\cdots+\int_{\mathcal{Y}}\sigma_{s}({\boldsymbol{x}},{\boldsymbol{y}})g_{s}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}}),$
where each $\sigma_{j}$, $j=0,1,\ldots,s$, are SOS polynomials. They also investigated the corresponding moment problem on the semialgebraic set $\mathcal{U}$. As an important application, their results could be used to solve semi-infinite optimization problems which are highly challenging.
Most Positivstellensätze could be generalized to the matrix setting that both $f$ and $g_{j}$’s are polynomial matrices. Scherer and Hol [42] developed a matrix-version of Putinar’s Positivstellensatz. Cimprič [6] extended the Krivine-Stengle Positivstellensatz to the case of polynomial matrices with polynomial constraints. A matrix version of Handelman’s Positivstellensatz was proposed in [27]. Building on Scherer and Hol’s Positivstellensatz, Dinh et al. [8] generalized the classical Schmüdgen, Putinar-Vasilescu, and Dickinson–Povh Positivstellensätze to the polynomial matrix setting. Given the projections of two semialgebraic sets defined by polynomial matrix inequalities (PMI), Klep and Nie [19] provided a matrix Positivstellensatz with lifting polynomials to determine whether one is contained in the other. These generalizations have a wide range of applications, particularly in areas such as optimal control, systems theory [13, 15, 17, 37, 47].
Computing SOS-structured representations involved in Positivstellensätze could be typically cast as semidefinite programs (SDP). However, the size of the SDPs grows rapidly with the dimension of the problem. Hence, from the perspective of computation, it becomes appealing to develop sparse versions of Positivstellensätze for sparse data, e.g., correlative sparsity [32, 23, 49], term sparsity [32, 29, 51, 52, 50], and matrix (chordal) sparsity [54, 53].
Due to estimation errors or lack of information, the data from real-world problems often involve uncertainty. As a result, ensuring the robustness of PMIs over a prescribed set with uncertainty is crucial for some safety-critical applications with little tolerance for failure [3]. Consequently, Positivstellensätze for polynomial matrices with UQs will be a powerful mathematical tool for addressing this issue, which serves as the primary motivation for this work. Precisely, consider the semialgebraic set
$\mathcal{X}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\subset{\mathbb{R}}^{m}\},$ (1)
where $G({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ (the set of $q\times q$ symmetric polynomial matrices in $\boldsymbol{x}$} and $\boldsymbol{y}$}), and $\mathcal{Y}\subset{\mathbb{R}}^{m}$ is closed. The goal of this paper is to provide certificates for positive definiteness of a $p\times p$ symmetric polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ over $\mathcal{X}$.
Table 1. Summary of Positivstellensätze for polynomials, polynomials with UQs, polynomial matrices, and polynomial matrices with UQs
Positivstellensatz Polynomials Polynomials with UQs PMIs PMIs with UQs
Putinar Putinar [39] Hu, Klep, and Nie [16] Scherer and Hol [42] Theorem 3.1
Putinar–Vasilescu Putinar and Vasilescu [38] Theorem 4.1 Dinh et al. [8] Theorem 4.1
Pólya Pólya [35] Theorem 4.2 Theorem 4.2 Theorem 4.2
Table 2. Summary of sparse Positivstellensätze for polynomials, polynomial matrices, and polynomial matrices with UQs
Case Literature
Polynomials with polynomial constraints Lasserre [23]
Polynomials with PMI constraints Kojima and Muramatsu [20]
Polynomial matrices with polynomial constraints Counterexample [32]
Polynomials with PMI constraints and UQs Theorem 3.2
Contributions. We generalize several classical Positivstellensätze from scalar polynomials to the matrix setting with UQs (see Tables 1 and 2 for summaries). In the following, we highlight the main results of this paper. Throughout the paper, let $\nu$ be a fixed Borel measure on ${\mathbb{R}}^{m}$ with support $\mathrm{supp}(\nu)=\mathcal{Y}$ and satisfying $\int_{\mathcal{Y}}|h({\boldsymbol{y}})|\mathrm{d}\nu({\boldsymbol{y}})<\infty$ for all $h({\boldsymbol{y}})\in{\mathbb{R}}[{\boldsymbol{y}}]$. For any $H({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{p}$, let us write $H({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\beta}}\in\mbox{supp}_{{\boldsymbol{y}}}(H)}H_{{\boldsymbol{\beta}}}({\boldsymbol{x}}){\boldsymbol{y}}^{{\boldsymbol{\beta}}}$ as a polynomial matrix in $\boldsymbol{y}$} with coefficient matrices $H_{{\boldsymbol{\beta}}}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and let
$\int_{\mathcal{Y}}H({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\coloneqq\sum_{{\boldsymbol{\beta}}\in\mbox{supp}_{{\boldsymbol{y}}}(H)}H_{{\boldsymbol{\beta}}}({\boldsymbol{x}})\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\beta}}}\mathrm{d}\nu({\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}.$
The main result of this paper is the following matrix-valued Positivestellensatz incorporating universal quantifiers, which holds under the Archimedean condition (Assumption 3.1) and the Carleman condition on $\nu$ (Assumption 3.2). See Section 2 for the definition of the product $\langle\cdot,\cdot\rangle_{p}$.
Theorem A. (Theorem 3.1) Suppose that Assumptions 3.1 and 3.2 hold. If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\mathcal{X}$, then there exist SOS matrices $_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ such that
$F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Then we consider the particular case of $p=1$, namely, the objective is a scalar polynomial. Suppose that $\mathcal{X}$ is defined by multiple PMIs with UQs ($G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,j=1,\ldots,s$). We assume the presence of correlative sparsity in the problem data, which implies that the variables $\boldsymbol{x}$} decompose as a union of subsets ${\boldsymbol{x}}=\cup_{\ell=1}^{t}{\boldsymbol{x}}(\mathcal{I}_{\ell})$ such that $\{G_{j}\}_{j=1}^{s}=\cup_{\ell=1}^{t}\{G_{j}\}_{j\in\mathcal{J}_{\ell}}$ and $G_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{q_{j}}$, $j\in\mathcal{J}_{\ell}$, $\ell=1,\ldots,t$ (see Assumption 3.3). Under these conditions, we could give the following sparse Positivstellensatz that leverages correlative sparsity patterns within these quantified constraints.
Theorem B. (Theorem 3.2) Suppose that $\mathcal{Y}$ is compact, Assumption 3.3 holds for $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ and $\mathcal{X}$, and Assumption 3.1 holds with respect to each ${\boldsymbol{x}}(\mathcal{I}_{\ell})$. If $f>0$ on $\mathcal{X}$, then there exist SOS polynomials $\sigma_{\ell,0}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$, and SOS polynomial matrices $S_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}]^{q_{j}}$, $j\in\mathcal{J}_{\ell}$, such that
$f({\boldsymbol{x}})=\sum_{\ell=1}^{t}\left(\sigma_{\ell,0}({\boldsymbol{x}}(\mathcal{I}_{\ell}))+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\langle S_{j}({\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}),G_{j}({\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}})\rangle\mathrm{d}\nu({\boldsymbol{y}})\right).$
Next, building on Theorem A and existing techniques, we establish two generalized Positivstellensätze for polynomial matrices with UQs and without assuming the Archimedean condition. The first is a Positivstellensatz for non-compact case.
Theorem C. (Theorem 4.1) Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$} of even degree, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $N\in\mathbb{N}$ and SOS matrices $_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in $\boldsymbol{x}$} and $\deg S_{0}=2N+\deg F$, $\deg_{{\boldsymbol{x}}}S=2N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
$\|{\boldsymbol{x}}\|^{2N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
The second is a Positivstellensatz on the nonnegative orthant.
Theorem D. (Theorem 4.2) Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$}, $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, and $\mathcal{Y}$ is compact. Then, there exists $N\in\mathbb{N}$ and polynomial matrices $P_{0}=\sum_{{\boldsymbol{\alpha}}}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $P=\sum_{{\boldsymbol{\alpha}}}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in $\boldsymbol{x}$} and satisfy each $P_{0,{\boldsymbol{\alpha}}}\succ 0$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all ${\boldsymbol{y}}\in\mathcal{Y}$, such that
$\left(\sum_{i=1}^{n}x_{i}\right)^{N}F({\boldsymbol{x}})=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
The inhomogeneous counterparts of Theorems C and D appear in Corollaries 4.1 and 4.3, respectively.
As applications, leveraging the SOS-structured certificates provided by the established Positivstellensätze, we propose hierarchies of SDP relaxations for the robust PMI constrained optimization problem and analyze their convergence.
The rest of this paper is organized as follows. In Section 2, we review some preliminary concepts. Section 3 presents a matrix-valued Positivstellensatz for polynomial matrices with UQs under the Archimedean condition, along with a sparse version in the presence of correlative sparsity. In Section 4, without relying on the Archimedean condition, we derive two generalized Positivstellensätze for polynomial matrices with UQs. In Section 5, we apply the established Positivstellensätze to robust PMI constrained optimization. For the sake of readability, some lengthy and technical proofs are deferred to Section 6. Conclusions are given in Section 7.
2. Preliminaries
[…]
$\mathrm{supp}(P)$ $\coloneqq\{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}\mid{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\ \text{appears in some }P_{ij}({\boldsymbol{x}})\},$
$\mbox{supp}_{{\boldsymbol{x}}}(Q)$ $\coloneqq\{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}\mid{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}{\boldsymbol{y}}^{{\boldsymbol{\beta}}}\ \text{appears in some }Q_{ij}({\boldsymbol{x}},{\boldsymbol{y}})\text{ for some }{\boldsymbol{\beta}}\in\mathbb{N}^{m}\},$
$\mbox{supp}_{{\boldsymbol{y}}}(Q)$ $\coloneqq\{{\boldsymbol{\beta}}\in\mathbb{N}^{m}\mid{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}{\boldsymbol{y}}^{{\boldsymbol{\beta}}}\ \text{appears in some }Q_{ij}({\boldsymbol{x}},{\boldsymbol{y}})\text{ for some }{\boldsymbol{\alpha}}\in\mathbb{N}^{n}\}.$
Let us recall some background on SOS matrices and positivstellensätz for polynomial matrices. For a polynomial $f({\boldsymbol{x}})\in{\mathbb{R}}[{\boldsymbol{x}}]$, if there exist polynomials $f_{1}({\boldsymbol{x}}),\ldots,f_{t}({\boldsymbol{x}})$ such that $f({\boldsymbol{x}})=\sum_{i=1}^{t}f_{i}({\boldsymbol{x}})^{2}$, then we call $f({\boldsymbol{x}})$ an SOS. A polynomial matrix $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is said to be an SOS matrix if there exists an $l\times p$ polynomial matrix $T({\boldsymbol{x}})$ for some $l\in\mathbb{N}$ such that $S({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})$. For $d\in\mathbb{N}$, denote by ${\boldsymbol{x}}]_{d}$ the canonical basis of ${\mathbb{R}}[{\boldsymbol{x}}]_{d}$, i.e.,
${\boldsymbol{x}}]_{d}\coloneqq[1,\ x_{1},\ x_{2},\ \cdots,\ x_{n},\ x_{1}^{2},\ x_{1}x_{2},\ \cdots,\ x_{n}^{d}]^{\intercal},$ (2)
whose cardinality is $|\mathbb{N}^{n}_{d}|=\binom{n+d}{d}$. With $d=\deg T$, we can write $T({\boldsymbol{x}})$ as
$T({\boldsymbol{x}})=Q([{\boldsymbol{x}}]_{d}\otimes I_{p})\text{ with }Q=[Q_{1},\ldots,Q_{|\mathbb{N}^{n}_{d}|}],\quad Q_{i}\in{\mathbb{R}}^{l\times p},$
where $Q$ is the vector of the coefficient matrices of $T({\boldsymbol{x}})$ with respect to ${\boldsymbol{x}}]_{d}$. Hence, $S({\boldsymbol{x}})$ is an SOS matrix with respect to ${\boldsymbol{x}}]_{d}$ if there exists some $Q\in{\mathbb{R}}^{l\times p|\mathbb{N}^{n}_{d}|}$ satisfying
$S({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})=([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}(Q^{\intercal}Q)([{\boldsymbol{x}}]_{d}\otimes I_{p}).$
Thus, we have the following results.
Proposition 2.1.
[42, Lemma 1] A polynomial matrix $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is an SOS matrix with respect to the monomial basis ${\boldsymbol{x}}]_{d}$ if and only if there exists $Z\in\mathbb{S}_{+}^{p|\mathbb{N}^{n}_{d}|}$ such that $S({\boldsymbol{x}})=([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}Z([{\boldsymbol{x}}]_{d}\otimes I_{p})$.
Lemma 2.1.
Let $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ be an SOS matrix and $S_{k}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{k\times k}$ be a principal submatrix of $S({\boldsymbol{x}})$ whose rows and columns are indexed by $(p_{1},\ldots,p_{k})$ with $1\leq p_{1}<\cdots<p_{k}\leq p$, then $S_{k}({\boldsymbol{x}})$ is an SOS matrix.
Proof.
As $S({\boldsymbol{x}})$ is an SOS matrix, there exists an $l\times p$ polynomial matrix $T({\boldsymbol{x}})$ for some $l\in\mathbb{N}$ such that $S({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})$. Denote by $T_{k}({\boldsymbol{x}})$ the submatrix of $T({\boldsymbol{x}})$ consisting of the columns of T(x)T(x)indexed by $(p_{1},\ldots,p_{k})$. Then $S_{k}({\boldsymbol{x}})=T_{k}({\boldsymbol{x}})^{\intercal}T_{k}({\boldsymbol{x}})$ and hence is an SOS matrix. ∎
We next recall Scherer-Hol’s Positivstellensatz for polynomial matrices obtained in [42]. Define the bilinear mapping
$\left\langle\cdot,\cdot\right\rangle_{p}\colon{\mathbb{R}}^{pq\times pq}\times{\mathbb{R}}^{q\times q}\to{\mathbb{R}}^{p\times p},\quad\left\langle A,B\right\rangle_{p}=\mbox{tr}_{p}\left(A^{\intercal}(I_{p}\otimes B)\right),$
with
$\mbox{tr}_{p}\left(C\right)\coloneqq\left[\begin{array}[]{ccc}\mbox{tr}\left(C_{11}\right)&\cdots&\mbox{tr}\left(C_{1p}\right)\\ \vdots&\ddots&\vdots\\ \mbox{tr}\left(C_{p1}\right)&\cdots&\mbox{tr}\left(C_{pp}\right)\end{array}\right]\quad\text{for }C=[C_{ij}]_{i,j\in[p]}\in{\mathbb{R}}^{pq\times pq},C_{ij}\in{\mathbb{R}}^{q\times q}.$
When $p=1$, the product $\left\langle A,B\right\rangle_{1}$ coincides with the matrix inner product $\langle A,B\rangle=\mbox{tr}\left(A^{\intercal}B\right)$.
Let $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$ where each $H_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$ for some $r_{j}\in\mathbb{N}$.
Lemma 2.2.
Let $H({\boldsymbol{x}})=\mbox{diag}(H_{1}({\boldsymbol{x}}),\ldots,H_{t}({\boldsymbol{x}}))$ be a block diagonal matrix, then for any SOS matrix $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{pr}$ where $p\in\mathbb{N}$ and $r=r_{1}+\cdots+r_{t}$, there are SOS matrices $S_{j}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$, $j\in[t]$, such that
$\left\langle S({\boldsymbol{x}}),H({\boldsymbol{x}})\right\rangle_{p}=\sum_{j=1}^{t}\left\langle S_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}.$
Proof.
For each $j\in[t]$, let $S_{j}({\boldsymbol{x}})$ be the $pr_{j}\times pr_{j}$ principal submatrix of $S({\boldsymbol{x}})$ whose rows and columns are indexed by
$\left(\sum_{\ell=1}^{j-1}r_{\ell}+1,\sum_{\ell=1}^{j}r_{\ell},r+\sum_{\ell=1}^{j-1}r_{\ell}+1,r+\sum_{\ell=1}^{j}r_{\ell},\ldots,(p-1)r+\sum_{\ell=1}^{j-1}r_{\ell}+1,(p-1)r+\sum_{\ell=1}^{j}r_{\ell}\right).$
As $H({\boldsymbol{x}})$ is block diagonal, by the definition of the mapping $\left\langle.,.\right\rangle_{p}$, it is easy to see that
$\left\langle S({\boldsymbol{x}}),H({\boldsymbol{x}})\right\rangle_{p}=\sum_{j=1}^{t}\left\langle S_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}.$
By Lemma 2.1, each $S_{j}({\boldsymbol{x}})$ is an SOS matrix. ∎
The matrix quadratic module $\mathcal{Q}^{p}(\mathbf{H})$ generated by $\mathbf{H}$ is defined as the cone
$\mathcal{Q}^{p}(\mathbf{H})\coloneqq\left\{S_{0}({\boldsymbol{x}})+\sum_{j=1}^{t}\left\langle S_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}\ \middle|\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{pr_{j}},j\in[t],\ \text{are SOS}\right\}.$
The $k$-th truncated matrix quadratic module $\mathcal{Q}^{p}_{k}(\mathbf{H})$ is a subcone of $\mathcal{Q}^{p}(\mathbf{H})$ in which the degrees of the polynomial matrices $S_{0}({\boldsymbol{x}})$ and $\sum_{j=1}^{t}\left\langle S_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}$ are bounded by $2k$.
Assumption 2.1.
$\mathcal{Q}^{p}(\mathbf{H})$ is Archimedean, i.e., there is $C>0$ such that $C-\|{\boldsymbol{x}}\|^{2}\in\mathcal{Q}^{1}(\mathbf{H})$.
Theorem 2.1.
(Scherer-Hol’s Positivestellensatz) Let Assumption 2.1 and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ be PD on $\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succeq 0,\ j\in[t]\}$. Then $F({\boldsymbol{x}})\in\mathcal{Q}^{p}(\mathbf{H})$.
Proof.
It was proved in [42, Corollary 1] for the case $t=1$. The case $t>1$ can be derived from [42, Corollary 1] and Lemma 2.2. ∎
3. Positivstellensätze for polynomial matrices with UQs under the Archimedean condition
In this section, assuming the Archimedean condition, we shall present a matrix-valued Positivstellensatz incorporating universal quantifiers, along with a sparse version in the presence of correlative sparsity.
3.1. A Positivstellensatze for polynomial matrices with UQs
Recall the set $\mathcal{X}$ in (1). Throughout the paper, let $\nu$ be a fixed Borel measure on ${\mathbb{R}}^{m}$ with support $\mathrm{supp}(\nu)=\mathcal{Y}$ and satisfying $\int_{\mathcal{Y}}|h({\boldsymbol{y}})|\mathrm{d}\nu({\boldsymbol{y}})<\infty$ for all $h({\boldsymbol{y}})\in{\mathbb{R}}[{\boldsymbol{y}}]$. Similarly to the scalar case in [16], let us define the matrix quadratic module associated with $(G,\nu)$ as follows.
Definition 3.1.
The matrix quadratic module $\mathcal{Q}^{p}(G,\nu)$ generated by $G$ and $\nu$ is defined as
$\mathcal{Q}^{p}(G,\nu)\coloneqq\left\{S_{0}+\int_{\mathcal{Y}}\left\langle S,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\ \middle|\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices}\right\}.$
Next, we derive a Positivstellensatze for polynomial matrices with UQs, providing SOS-structured characterizations for polynomial matrices that are PD over $\mathcal{X}$.
Assumption 3.1.
$\mathcal{Q}^{p}(G,\nu)$ is Archimedean, i.e., there is $C>0$ such that $C-\|{\boldsymbol{x}}\|^{2}\in\mathcal{Q}^{1}(G,\nu)$.
Consider the Carleman condition imposed on $\nu$ which is automatically satisfied when $\mathcal{Y}$ is compact.
Assumption 3.2.
The Borel measure $\nu$ satisfies the multivariate Carleman condition
$\sum_{d=1}^{\infty}\left(\int_{\mathcal{Y}}y_{j}^{2d}\mathrm{d}\nu({\boldsymbol{y}})\right)^{-\frac{1}{2d}}=\infty,\quad\forall j\in[m].$
Proposition 3.1.
[16, Proposition 3.2] Suppose that $\nu$ satisfies Assumption 3.2. Then, SOS polynomials are dense in the cone of nonnegative functions in $L^{2}({\mathbb{R}}^{m},\nu)$.
Proposition 3.2.
Suppose that $\nu$ satisfies Assumption 3.2. Then
$\mathcal{X}=\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0,\ \forall H\in\mathcal{Q}^{p}(G,\nu)\}.$ (3)
Proof.
By the definition of $\mathcal{Q}^{p}(G,\nu)$, we only need to prove that $\mathcal{X}$ contains the set on the right-hand side of the equation in (3). Fix ${\boldsymbol{u}}\in{\mathbb{R}}^{n}$ with $H({\boldsymbol{u}})\succeq 0$ for all $H\in\mathcal{Q}^{p}(G,\nu)$. Suppose on the contrary that ${\boldsymbol{u}}\not\in\mathcal{X}$, i.e., there exists ${\boldsymbol{w}}\in\mathcal{Y}$ such that $G({\boldsymbol{u}},{\boldsymbol{w}})\not\succeq 0$. Then, there is a ball $\mathcal{O}\subset{\mathbb{R}}^{m}$ with a radius $\rho>0$ around $\boldsymbol{w}$} such that $G({\boldsymbol{u}},{\boldsymbol{y}})\not\succeq 0$ on $2\mathcal{O}$. We may assume that there exists ${\boldsymbol{v}}\in{\mathbb{R}}^{q}$ and $\delta>0$ such that ${\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}\leq-\delta$ on $2\mathcal{O}$. Define a continuous function $h({\boldsymbol{y}})$ on ${\mathbb{R}}^{m}$ by $h({\boldsymbol{y}})=2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|$ for ${\boldsymbol{y}}\in 2\mathcal{O}$ and $h({\boldsymbol{y}})=0$ otherwise. By Proposition 3.1, there exists a sequence of SOS polynomials $\{\sigma_{k}\}_{k}$ in ${\mathbb{R}}[{\boldsymbol{y}}]$ that converges to $h$ in the $L^{2}$-norm. Hence,
$\lim_{k\to\infty}\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})=\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}h({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$
$=$ $\int_{\mathcal{Y}\cap 2\mathcal{O}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}(2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|)\mathrm{d}\nu({\boldsymbol{y}})\leq\int_{\mathcal{Y}\cap\mathcal{O}}-\delta\rho\mathrm{d}\nu({\boldsymbol{y}})=-\delta\rho\nu(\mathcal{Y}\cap\mathcal{O})<0,$
where the last inequality is due to the fact that $\mathrm{supp}(\nu)=\mathcal{Y}$ and thus $\nu(\mathcal{Y}\cap\mathcal{O})>0$. Note that for all $k\in\mathbb{N}$,
$\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})=\frac{1}{p}\mbox{tr}\left(\int_{\mathcal{Y}}\left\langle\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal},G({\boldsymbol{u}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\geq 0,$
since $\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}$ is an SOS and $H({\boldsymbol{u}})\succeq 0$ for all $H\in\mathcal{Q}^{p}(G,\nu)$. A contradiction follows. ∎
Now, we present our main result concerning a Positivstellensatz for polynomial matrices with UQs.
Theorem 3.1.
Suppose that Assumptions 3.1 and 3.2 hold. If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\mathcal{X}$, then $F({\boldsymbol{x}})\in\mathcal{Q}^{p}(G,\nu)$.
Proof.
Since the quadratic module $\mathcal{Q}^{p}(G,\nu)$ is Archimedean and the equality in (3) holds, the conclusion follows from the fundamental Positivstellensatz for matrix algebras of polynomials [44, Theorem 10.25]. ∎
Next, we derive a corollary of Theorem 3.1, which will be used in Section 4.
Lemma 3.1.
Suppose that $S({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ is an SOS matrix in $\boldsymbol{x}$} and $\boldsymbol{y}$}. Then $\int_{\mathcal{Y}}S({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is an SOS matrix in $\boldsymbol{x}$}.
Proof.
As $S({\boldsymbol{x}},{\boldsymbol{y}})$ is an SOS matrix, there exists an $\ell\times q$ polynomial matrix $T({\boldsymbol{x}},{\boldsymbol{y}})$ for some $\ell\in\mathbb{N}$ such that $S({\boldsymbol{x}},{\boldsymbol{y}})=T({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}T({\boldsymbol{x}},{\boldsymbol{y}})$. With $d=\deg_{{\boldsymbol{x}}}T$, we could write $T({\boldsymbol{x}},{\boldsymbol{y}})$ as
$T({\boldsymbol{x}},{\boldsymbol{y}})=Q({\boldsymbol{y}})([{\boldsymbol{x}}]_{d}\otimes I_{q})\text{ with }Q({\boldsymbol{y}})=[Q_{1}({\boldsymbol{y}}),\ldots,Q_{|\mathbb{N}^{n}_{d}|}({\boldsymbol{y}})],\quad Q_{i}({\boldsymbol{y}})\in{\mathbb{R}}[{\boldsymbol{y}}]^{\ell\times q},$
where $Q({\boldsymbol{y}})$ is the vector of coefficient matrices of $T({\boldsymbol{x}},{\boldsymbol{y}})$ (considered as a polynomial matrix in ${\mathbb{R}}[{\boldsymbol{x}}]^{\ell\times q}$) with respect to ${\boldsymbol{x}}]_{d}$. Hence,
$\int_{\mathcal{Y}}S({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})=\int_{\mathcal{Y}}([{\boldsymbol{x}}]_{d}\otimes I_{q})^{\intercal}(Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}}))([{\boldsymbol{x}}]_{d}\otimes I_{q})\mathrm{d}\nu({\boldsymbol{y}})=([{\boldsymbol{x}}]_{d}\otimes I_{q})^{\intercal}\int_{\mathcal{Y}}Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})([{\boldsymbol{x}}]_{d}\otimes I_{q})$
As $\int_{\mathcal{Y}}Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is PSD, $\int_{\mathcal{Y}}S({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is an SOS matrix in $\boldsymbol{x}$}. ∎
Corollary 3.1.
Let $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$ where each $H_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$ for some $r_{j}\in\mathbb{N}$. Suppose that Assumption 3.2 holds and there exists $C>0$ such that $C-\|{\boldsymbol{x}}\|^{2}\in\mathcal{Q}^{1}(G,\nu)+\mathcal{Q}^{1}(\mathbf{H})$. If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succeq 0,\ j\in[t]\}$, then $F({\boldsymbol{x}})\in\mathcal{Q}^{p}(G,\nu)+\mathcal{Q}^{p}(\mathbf{H})$.
Proof.
It is clear that $F({\boldsymbol{x}})\succ 0$ on $\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succeq 0,\ j\in[t]\}$ if and only if $F({\boldsymbol{x}})\succ 0$ on
$\left\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\,\middle|\,\widehat{G}({\boldsymbol{x}},{\boldsymbol{y}})\coloneqq\mbox{diag}(G({\boldsymbol{x}},{\boldsymbol{y}}),H_{1}({\boldsymbol{x}}),\ldots,H_{t}({\boldsymbol{x}}))\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\right\}.$
By Lemmas 2.2 and 3.1, it is easy to see that $\mathcal{Q}^{p}(\widehat{G},\nu)=\mathcal{Q}^{p}(G,\nu)+\mathcal{Q}^{p}(\mathbf{H})$. So, the conclusion follows from Theorem 3.1. ∎
3.2. A sparse Positivstellensatz for polynomial matrices with UQs
There are sparse Positivstellensätze for scalar polynomials being positive on a basic semialgebraic set in the presence of correlative sparsity; see [11, 23] where the set is defined by polynomial inequalities and [20] where the set is defined PMIs. Let $p=1$. Building on the ideas from [11] and [20], we now prove a sparse Positivstellensatz for polynomial matrices with UQs. Let $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ and
$\widehat{\mathcal{X}}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall j\in[s],\ {\boldsymbol{y}}\in\mathcal{Y}\subset{\mathbb{R}}^{m}\},$ (4)
where each $G_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q_{j}}$, $q_{j}\in\mathbb{N}$.
Assumption 3.3 (correlative sparsity pattern).
Subsets $\{\mathcal{I}_{\ell}\}_{\ell\in[t]}$ of $[n]$ and subsets $\{\mathcal{J}_{\ell}\}_{\ell\in[t]}$ of $[s]$ satisfy the following conditions:
  1. (i)
    The running intersection property holds for $\{\mathcal{I}_{\ell}\}_{\ell\in[t]}$, i.e.,
    $\text{for }\ell=2,\ldots,t,\ \exists k<\ell\quad\text{s.t.}\quad\mathcal{I}_{\ell}\cap\bigcup_{j<\ell}\mathcal{I}_{j}\subseteq\mathcal{I}_{k};$
  2. (ii)
    For every $\ell\in[t]$ and $j\in\mathcal{J}_{\ell}$, $G_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}]^{q_{j}}$, where ${\boldsymbol{x}}(\mathcal{I}_{\ell})\coloneqq\{x_{i}\}_{i\in\mathcal{I}_{\ell}}$;
  3. (iii)
    $f$ decomposes as $f=f_{1}+\cdots+f_{t}$ with each $f_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$.
For each $\ell\in[t]$, let $\mathbf{G}^{\ell}\coloneqq\{G_{j}\}_{j\in\mathcal{J}_{\ell}}$ and $\mathcal{Q}^{p}(\mathbf{G}^{\ell},\nu)$ be the quadratic module generated by $\mathbf{G}^{\ell}$ and $\nu$ in ${\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{p}$, i.e.,
$\mathcal{Q}^{p}(\mathbf{G}^{\ell},\nu)\coloneqq\left\{S_{0}+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\left\langle S_{j},G_{j}\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\ \middle|\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{p},S_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}]^{pq_{j}},j\in\mathcal{J}_{\ell},\text{ are SOS matrices}\right\}.$
Using the correlative sparsity pattern and the Archimedean condition, we are able to derive the following sparse Positivstellensatz.
Theorem 3.2.
Suppose that $\mathcal{Y}$ is compact, Assumption 3.3 holds for $f$ and $\widehat{\mathcal{X}}$, and Assumption 3.1 holds for each $\mathcal{Q}^{1}(\mathbf{G}^{\ell},\nu)$, $\ell\in[t]$. If $f>0$ on $\widehat{\mathcal{X}}$, then $f\in\sum_{\ell=1}^{t}\mathcal{Q}^{1}(\mathbf{G}^{\ell},\nu)$.
To prove Theorem 3.2, we need the following intermediary results.
Proposition 3.3.
[11, Lemma 3] Let $\{\mathcal{I}_{\ell}\}_{\ell\in[t]}$ satisfy the running intersection property. For any $C>0$, if $f=f_{1}+\cdots+f_{t}$ with $f_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ satisfies $f>0$ on $C,C]^{n}$, then $f=h_{1}+\ldots+h_{t}$ for some $h_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ with $h_{\ell}>0$ on $C,C]^{|\mathcal{I}_{\ell}|}$.
Proposition 3.4.
Suppose $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\widehat{\mathcal{X}}$. Then for any $C>0$, there exists $M>0$ and $\bar{k}\in\mathbb{N}$ such that
$F({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\succ 0$ (5)
on $C,C]^{n}$ for all $k\geq\bar{k}$.
The proof of Proposition 3.4 is postponed to Section 6.1. Note that using Proposition 3.4, we can provide a constructive proof of Theorem 3.1 assuming that $\mathcal{Y}$ is compact (see Section 6.1).
Proof of Theorem 3.2.
By assumption, let $C>0$ be such that $C-\|{\boldsymbol{x}}(\mathcal{I}_{\ell})\|^{2}\in\mathcal{Q}^{1}(\mathbf{G}^{\ell},\nu)$ holds for all $\ell\in[t]$. By Proposition 3.4, there exists $M>0$ and $k^{\prime}\in\mathbb{N}$ such that
$f({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k^{\prime}},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})>0$
on $\sqrt{C},\sqrt{C}]^{n}$. Note that for each $j\in[s]$,
$\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k^{\prime}},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$
for some $\ell\in[t]$. By Proposition 3.3, there exist $h_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$, $\ell\in[t]$, such that
$f({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k^{\prime}},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})=h_{1}+\cdots+h_{t},$
and each $h_{\ell}>0$ on $\sqrt{C},\sqrt{C}]^{|\mathcal{I}_{\ell}|}$. By Putinar’s Positivstellensatz [39], for each $h_{\ell}$, there exist SOS polynomials $\sigma_{\ell,0},\sigma_{\ell,1}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ such that
$h_{\ell}=\sigma_{\ell,0}+\sigma_{\ell,1}(C-\|{\boldsymbol{x}}(\mathcal{I}_{\ell})\|^{2}).$
As each $C-\|{\boldsymbol{x}}(\mathcal{I}_{\ell})\|^{2}\in\mathcal{Q}^{1}(\mathbf{G}^{\ell},\nu)$, it holds $h_{\ell}\in\mathcal{Q}^{1}(\mathbf{G}^{\ell},\nu)$. Therefore,
$f({\boldsymbol{x}})=\sum_{\ell\in[t]}\left(\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k^{\prime}},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})+h_{\ell}\right)\in\sum_{\ell=1}^{t}\mathcal{Q}^{1}(\mathbf{G}^{\ell},\nu).$
Remark 3.1.
One might wonder whether the result of Theorem 3.2 holds for a polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ with $p>1$. Indeed, this is not true even in the absence of UQs; see [32] for a counterexample.
4. Positivstellensätze for polynomial matrices with UQs and without the Archimedean condition
The results in Section 3 are derived under the Archimedean condition on the quadratic module generated by $(G,\nu)$, which requires $\mathcal{X}$ to be bounded. Many efforts have been made to provide SOS-structured representations for polynomials that are PD on a semialgebraic set without the Archimedean condition. The goal of this section is to extend some of these well-known results to the setting of polynomial matrices with UQs.
4.1. A Positivstellensatz for the non-compact case
Let $\theta\coloneqq 1+\|{\boldsymbol{x}}\|^{2}$. For every nonnegative polynomial $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ on a general basic semialgebraic set, Putinar and Vasilescu [38] proved that for a given $\varepsilon>0$ and $2d\geq\deg f$, there exists a nonnegative integer $k$ such that $\theta^{k}(f+\varepsilon\theta^{d})$ belongs to the quadratic module associated with the basic semialgebraic set. Using Jacobi’s technique [18], Mai et al. [30] provided an alternative proof of Putinar and Vasilescu’s result with an effective degree bound on polynomials involved in such certificates.
Building on Theorem 3.1 and similar techniques from [30], we next derive matrix-valued Positivstellensätz for the non-compact case, incorporating universal quantifiers. In the homogeneous case, the result is stated as follows. See Section 6.2 for its proof.
Theorem 4.1.
Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$} of even degree, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $N\in\mathbb{N}$ and SOS matrices $_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ that are homogeneous in $\boldsymbol{x}$} and $\deg S_{0}=2N+\deg F$, $\deg_{{\boldsymbol{x}}}S=2N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
$\|{\boldsymbol{x}}\|^{2N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
For any $H=[H_{ij}]\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{r}$, let $d_{H}\coloneqq\max\left\{\lfloor\deg_{{\boldsymbol{x}}}(H_{ij})/2\rfloor+1:i,j\in[r]\right\}$ and $\tilde{{\boldsymbol{x}}}\coloneqq({\boldsymbol{x}},x_{n+1})$. By applying Theorem 4.1, we obtain its inhomogeneous counterpart, as given below.
Corollary 4.1.
Suppose that Assumption 3.2 holds and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}$. Then for any $\varepsilon>0$, there exists $N_{\varepsilon}\in\mathbb{N}$, SOS matrices $S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, and $S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg S_{0}\leq 2(N_{\varepsilon}+d_{F})$ and $\deg_{{\boldsymbol{x}}}S\leq 2(N_{\varepsilon}+d_{F}-d_{G})$, such that
$\theta^{N_{\varepsilon}}(F({\boldsymbol{x}})+\varepsilon\theta^{d_{F}}I_{p})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Proof.
Let $\widetilde{F}\coloneqq[x_{n+1}^{2d_{F}}F_{ij}({\boldsymbol{x}}/x_{n+1})]_{i,j\in[p]}\in{\mathbb{S}}[\tilde{{\boldsymbol{x}}}]^{p}$, $\widetilde{G}\coloneqq[x_{n+1}^{2d_{G}}G_{ij}({\boldsymbol{x}}/x_{n+1},{\boldsymbol{y}})]_{i,j\in[q]}\in{\mathbb{S}}[\tilde{{\boldsymbol{x}}},{\boldsymbol{y}}]^{q}$, and consider
$\widetilde{\mathcal{X}}\coloneqq\left\{\tilde{{\boldsymbol{x}}}\in{\mathbb{R}}^{n+1}\middle|\widetilde{G}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}})\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\right\}.$
We first prove that $\widetilde{F}+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p}\succ 0$ on $\widetilde{\mathcal{X}}\setminus\{\mathbf{0}\}$. Fix a point $\tilde{{\boldsymbol{u}}}=({\boldsymbol{u}},u_{n+1})\in\widetilde{\mathcal{X}}\setminus\{\mathbf{0}\}$.
  1. -
    Case 1: $u_{n+1}\neq 0$. As $\widetilde{G}(\tilde{{\boldsymbol{u}}},{\boldsymbol{y}})=u_{n+1}^{2d_{G}}G({\boldsymbol{u}}/u_{n+1},{\boldsymbol{y}})\succeq 0$, we have $G({\boldsymbol{u}}/u_{n+1},{\boldsymbol{y}})\succeq 0$ for all ${\boldsymbol{y}}\in\mathcal{Y}$, which implies ${\boldsymbol{u}}/u_{n+1}\in\mathcal{X}$. Hence, $\widetilde{F}(\tilde{{\boldsymbol{u}}})=u_{n+1}^{2d_{F}}F({\boldsymbol{u}}/x_{n+1})\succeq 0$. Since $\|\tilde{{\boldsymbol{u}}}\|^{2}\neq 0$, $\widetilde{F}(\tilde{{\boldsymbol{u}}})+\varepsilon\|\tilde{{\boldsymbol{u}}}\|^{2d_{F}}I_{p}\succ 0$.
  2. -
[…]
Applying Theorem 4.1 to the polynomial matrices $\tilde{F}+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p}$ and $\widetilde{G}$, we obtain $N_{\varepsilon}\in\mathbb{N}$ and SOS matrices $\widetilde{S}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\widetilde{S}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$, which are homogeneous in $\boldsymbol{x}$} and $\deg\widetilde{S}_{0}=2N_{\varepsilon}+2d_{F}$, $\deg_{{\boldsymbol{x}}}\widetilde{S}=2N_{\varepsilon}+2d_{F}-2d_{G}$, such that
$\|\tilde{{\boldsymbol{x}}}\|^{2N_{\varepsilon}}(\widetilde{F}(\tilde{{\boldsymbol{x}}})+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p})=\widetilde{S}_{0}(\tilde{{\boldsymbol{x}}})+\int_{\mathcal{Y}}\left\langle\widetilde{S}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}}),\widetilde{G}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$ (6)
Then, letting $x_{n+1}=1$ in (6), we achieve the desired conclusion. ∎
4.2. A Positivstellensatz on the nonnegative orthant
Pólya [35] proved that multiplying a positive homogeneous polynomial on the nonnegative orthant by some power of $\sum_{i=1}^{n}x_{i}$ yields a polynomial with nonnegative coefficients. Powers and Reznick [36] gave a ”positive” version of Pólya’s theorem. Dickinson and Povh [7] extended the result of Pólya [35] to provide a certificate for positive homogeneous polynomials on the intersection of the nonnegative orthant with a basic semialgebraic set.
We next employ Theorem 4.1 and similar techniques from [7] to present a Positivstellensatz for homogeneous polynomial matrices being PD on the intersection of $\mathcal{X}$ with the nonnegative orthant. Denote by $\boldsymbol{e}$} the column vector in ${\mathbb{R}}^{n}$ of all ones.
Theorem 4.2.
Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$}, $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, and $\mathcal{Y}$ is compact. Then, there exists $N\in\mathbb{N}$ and polynomial matrices
$P_{0}=\sum_{|{\boldsymbol{\alpha}}|=N+\deg F}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\quad P=\sum_{|{\boldsymbol{\alpha}}|=N+\deg F-\deg_{{\boldsymbol{x}}}G}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$
which are homogeneous in $\boldsymbol{x}$} and satisfy $P_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=N+\deg F$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all $|{\boldsymbol{\alpha}}|=N+\deg F-\deg_{{\boldsymbol{x}}}G$ and ${\boldsymbol{y}}\in\mathcal{Y}$, such that
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
The proof is postponed to Section 6.3.
In particular, when $\mathcal{Y}$ is a semialgebriac set defined by PMIs, the coefficient matrices $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})$ in Theorem 4.2 have SOS-structured representations.
Corollary 4.2.
Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$}, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$. Moreover, suppose that
$\mathcal{Y}=\{{\boldsymbol{y}}\in{\mathbb{R}}^{m}\mid H_{1}({\boldsymbol{y}})\succeq 0,\ldots,H_{t}({\boldsymbol{y}})\succeq 0\},\ H_{i}({\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{y}}]^{\ell_{i}},\ \ell_{i}\in\mathbb{N},\ i\in[t],$ (7)
and the Archimedean condition holds for the quadratic module $\mathcal{Q}^{pq}(\mathbf{H})$ where $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$. Then, there exists $N\in\mathbb{N}$ and polynomial matrices
$P_{0}=\sum_{|{\boldsymbol{\alpha}}|=N+\deg F}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\quad P=\sum_{|{\boldsymbol{\alpha}}|=N+\deg F-\deg_{{\boldsymbol{x}}}G}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$
which are homogeneous in $\boldsymbol{x}$} and satisfy $P_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=N+\deg F$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\in\mathcal{Q}^{pq}(\mathbf{H})$ for all $|{\boldsymbol{\alpha}}|=N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Proof.
It follows from Theorem 4.2 and Scherer-Hol’s Positivestellensatz (Theorem 2.1). ∎
In analogy with Corollary 4.1, we can obtain the following inhomogeneous counterpart of Theorem 4.2. We omit the similar proof for the sake of clarity.
Corollary 4.3.
Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$, $F({\boldsymbol{x}})$ is PSD on ${\mathbb{R}}^{n}_{+}\cap\mathcal{X}$, and $\mathcal{Y}$ is compact. Then for any $\varepsilon>0$, there exists $N_{\varepsilon}\in\mathbb{N}$ and polynomial matrices
$P_{0}=\sum_{|{\boldsymbol{\alpha}}|\leq N_{\varepsilon}+2d_{F}}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\quad P=\sum_{|{\boldsymbol{\alpha}}|\leq N_{\varepsilon}+2d_{F}-2d_{G}}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$
satisfying $P_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|\leq N_{\varepsilon}+2d_{F}$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all $|{\boldsymbol{\alpha}}|\leq N_{\varepsilon}+2d_{F}-2d_{G}$ and ${\boldsymbol{y}}\in\mathcal{Y}$, and
$(1+{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{\varepsilon}}\left(F({\boldsymbol{x}})+\varepsilon(1+{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{2d_{F}}I_{p}\right)=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
5. Applications to robust PMI constrained optimization
Verifying PMIs over a prescribed set has a wide range of applications in many fields. For example, many control problems for systems of ordinary differential equations can be formulated as convex optimization problems with PMI constraints that must be satisfied over a specified portion of the state space [13, 15, 17, 37, 47]. These problems are typically formulated as follows:
$\inf_{{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}}\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\ \text{ s.t. }\ \Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})\coloneqq\Lambda_{0}({\boldsymbol{x}})-\sum_{i=1}^{r}\Lambda_{i}({\boldsymbol{x}})\gamma_{i}\succeq 0,\ \forall{\boldsymbol{x}}\in\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0\},$
where ${\boldsymbol{c}}\in{\mathbb{R}}^{r}$, $\Lambda_{0},\ldots,\Lambda_{r}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $H\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$.
However, due to estimation errors or lack of information, data from real-world problems often involve uncertainty. Therefore, ensuring the robustness of PMIs over a prescribed set under uncertainty is a critical issue, which could be formulated as the following robust optimization problem:
$\tau^{\star}\coloneqq\inf_{{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}}\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\ \text{ s.t. }\ \Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})\succeq 0,\ \forall{\boldsymbol{x}}\in\mathcal{X}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\},$ (8)
where $G({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ and $\mathcal{Y}\subset{\mathbb{R}}^{m}$ is closed.
Consequently, Positivstellensätze for polynomial matrices with UQs developed in this paper serve as a powerful mathematical tool for addressing this issue. Indeed, by leveraging the SOS-structured certificates provided by Positivstellensätze for the positive definiteness of $\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})$ over $\mathcal{X}$, we are able to establish hierarchies of SDP relaxations for the problem (8).
5.1. The compact case
If $\mathcal{X}$ is compact, then one could apply Theorem 3.1 to construct a hierarchy of SDP relaxations for (8):
$\left\{\begin{aligned} \tau(k)\coloneqq\inf_{{\boldsymbol{\gamma}},S_{0},S}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ \Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\ \deg S_{0}\leq 2k,\ \deg_{{\boldsymbol{x}}}S\leq 2(k-\lceil\deg_{{\boldsymbol{x}}}G/2\rceil),\ \deg_{{\boldsymbol{y}}}S\leq 2(k-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil).\end{aligned}\right$. (9)
It is clear that $\{\tau(k)\}_{k\in\mathbb{N}}$ is a sequence of non-increasing upper bounds on $\tau^{\star}$.
Theorem 5.1.
Suppose that Assumptions 3.1 and 3.2 hold and there exists a point $\bar{{\boldsymbol{\gamma}}}\in{\mathbb{R}}^{r}$ such that $\Lambda({\boldsymbol{x}},\bar{{\boldsymbol{\gamma}}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}$. Then we have $\lim_{k\to\infty}\tau(k)=\tau^{\star}$.
Proof.
Note that for any feasible point $\boldsymbol{\gamma}$$^{\prime}$ of (8) and any $\varepsilon>0$, the point $(1-\varepsilon){\boldsymbol{\gamma}}^{\prime}+\varepsilon\bar{{\boldsymbol{\gamma}}}$ satisfies $\Lambda({\boldsymbol{x}},(1-\varepsilon){\boldsymbol{\gamma}}^{\prime}+\varepsilon\bar{{\boldsymbol{\gamma}}})\succ 0$ on $\mathcal{X}$. Then, the conclusion follows by Theorem 3.1. ∎
Given that the SDP problem derived from (9) scales rapidly and becomes computationally prohibitive as $k$ increases, we revise it by imposing separate degree constraints on $\boldsymbol{x}$} and $\boldsymbol{y}$}, leading to the following problem
$\left\{\begin{aligned} \tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})\coloneqq\inf_{{\boldsymbol{\gamma}},S_{0},S}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ \Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\ \deg S_{0}\leq 2k_{{\boldsymbol{x}}},\ \deg_{{\boldsymbol{x}}}S\leq 2(k_{{\boldsymbol{x}}}-\lceil\deg_{{\boldsymbol{x}}}G/2\rceil),\ \deg_{{\boldsymbol{y}}}S\leq 2(k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil).\end{aligned}\right$. (10)
If $\mathcal{Y}$ is low-dimensional, we can maintain a small $k_{{\boldsymbol{x}}}$ while increasing $k_{{\boldsymbol{y}}}$ separately; this allows us to solve (10) in reasonable time and obtain a nontrivial upper bound $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ of $\tau^{\star}$.
In practice, the SDP relaxation (10) requires an efficient evaluation of the integral $\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\beta}}}\mathrm{d}\nu({\boldsymbol{y}})$. If $\nu$ is chosen to be the Lebesgue measure on $\mathcal{Y}$ (as in the following illustrating examples), this integral is available in closed form or by numerical methods whenever $\mathcal{Y}$ is a simple domain; examples include hypercubes, spheres [9], balls [9], and polytopes [2]. All numerical experiments in the sequel were carried out on a PC with an 8-Core Intel i7 3.8GHz CPU and 8G RAM, using the SDP solver MOSEK [1] and a customized version of the MATLAB optimization toolbox YALMIP [28].
Example 5.1.
Consider the feasible set of the problem (8):
$\mathcal{F}=\{{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}\mid\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})\succeq 0,\forall{\boldsymbol{x}}\in\mathcal{X}\}.$
In this example, we illustrate how Theorem 3.1 enables one to approximate the convex set of vectors $\boldsymbol{\gamma}$} for which $\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})$ is PSD on $\mathcal{X}$ – that is, to obtain approximate semidefinite representations of $\mathcal{F}$.
For any $k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}\in\mathbb{N}$ with $k_{{\boldsymbol{x}}}\geq\lceil\deg_{{\boldsymbol{x}}}G/2\rceil$ and $k_{{\boldsymbol{y}}}\geq\lceil\deg_{{\boldsymbol{y}}}G/2\rceil$, let $\nu$ in (10) be the Lebesgue measure on $\mathcal{Y}$ and consider
$\mathcal{F}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})=\left\{\begin{array}[]{c|c}{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}&\begin{aligned} &\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}{\boldsymbol{y}},\\ &S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\deg S_{0}\leq 2k_{{\boldsymbol{x}}},\ \deg_{{\boldsymbol{x}}}S\leq 2(k_{{\boldsymbol{x}}}-\lceil\deg_{{\boldsymbol{x}}}G/2\rceil),\ \deg_{{\boldsymbol{y}}}S\leq 2(k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil).\end{aligned}\end{array}\right\}.$
Clearly, $\mathcal{F}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ is an inner approximation of $\mathcal{F}$. On the other hand, Theorem 3.1 guarantees that any ${\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}$ for which $\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})$ is PD on $\mathcal{X}$ belongs to $\mathcal{F}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ for sufficiently large $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$. Therefore, the sets $\mathcal{F}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ can approximate $\mathcal{F}$ from inside arbitrarily accurately in the sense that any compact subset of the interior of $\mathcal{F}$ is included in $\mathcal{F}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ for some sufficiently large integers $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$.
Let us construct an illustrating example. Let $n=r=2$, $m=1$, $\mathcal{Y}=[0,1]$,
$G({\boldsymbol{x}},y)=\left[\begin{array}[]{cc}(1/2-y)^{2}&y(x_{1}^{2}+x_{2}^{2}-1)\\ y(x_{1}^{2}+x_{2}^{2}-1)&x_{1}+x_{2}-1\end{array}\right],$
and
$\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})=\left[\begin{array}[]{cc}1-(x_{1}\gamma_{1}-x_{2}\gamma_{2}+1/2)&2(x_{1}\gamma_{2}+x_{2}\gamma_{1})\\ 2(x_{1}\gamma_{2}+x_{2}\gamma_{1})&1+(x_{1}\gamma_{1}-x_{2}\gamma_{2}+1/2)\end{array}\right].$
It is clear that
$\mathcal{X}=\{{\boldsymbol{x}}\in{\mathbb{R}}^{2}\mid x_{1}^{2}+x_{2}^{2}=1,\ x_{1}x_{2}\geq 0\}.$
Geometrically, the set $\mathcal{F}$ is constructed as the intersection of all rotated copies of the shape defined by $(\gamma_{1}+1/2)^{2}+4\gamma_{2}^{2}\leq 1$ in the $\boldsymbol{\gamma}$}-plane as it is continuously rotated clockwise by $90^{\circ}$ about the origin. The resulting set $\mathcal{F}$ (light blue region with blue outline) is shown in Figure 1. Three inner approximations – $\mathcal{F}(1,1)$, $\mathcal{F}(1,2)$, and $\mathcal{F}(1,6)$ – each depicted with a red outline, appear in Figures 1. These red outlines were obtained by solving and extracting optimal solutions of (8) with objective functions $\gamma_{1}\cos\theta+\gamma_{2}\sin\theta$ over $80$ equispaced values of $\theta$ in the interval $0,2\pi]$.
Figure 1. The region $\mathcal{F}$ and three inner approximations $\mathcal{F}(1,1)$, $\mathcal{F}(1,2)$, $\mathcal{F}(1,6)$ in Example 5.1.
5.2. Exploiting sparsity
We illustrate the computational advantage of our sparsity-exploiting Positivstellensatz when the correlative sparsity pattern in Assumption 3.3 is satisfied. Consider the problem
$f^{\star}\coloneqq\min\ f({\boldsymbol{x}})\ \text{ s.t. }{\boldsymbol{x}}\in\widehat{\mathcal{X}}$,
where $f({\boldsymbol{x}})\in{\mathbb{R}}[{\boldsymbol{x}}]$ and the set $\widehat{\mathcal{X}}$ is defined in (4). By Theorem 3.1, dense SDP relaxations for such a problem can be formulated as
$\left\{\begin{aligned} \tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})\coloneqq\sup_{\gamma,\sigma_{0},S_{j}}&\ \gamma\\ \text{s.t.}&\ f({\boldsymbol{x}})-\gamma=\sigma_{0}({\boldsymbol{x}})+\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle S_{j}({\boldsymbol{x}},{\boldsymbol{y}}),G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ \sigma_{0}\in{\mathbb{R}}[{\boldsymbol{x}}],S_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q_{j}},\,j\in[s],\,\ \ \text{are SOS},\\ &\ \deg\sigma_{0}\leq 2k_{{\boldsymbol{x}}},\ \deg_{{\boldsymbol{x}}}S_{j}\leq 2(k_{{\boldsymbol{x}}}-\lceil\deg_{{\boldsymbol{x}}}G_{j}/2\rceil),\ \deg_{{\boldsymbol{y}}}S_{j}\leq 2(k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G_{j}/2\rceil).\end{aligned}\right$. (11)
For any $k_{{\boldsymbol{x}}}\geq\lceil\deg_{{\boldsymbol{x}}}G_{j}/2\rceil$ and $k_{{\boldsymbol{y}}}\geq\lceil\deg_{{\boldsymbol{y}}}G_{j}/2\rceil$, the number $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ provides a lower bound of $f^{\star}$. The sequence $\{\tau(k,k)\}_{k\in\mathbb{N}}$, the limit $\tau(\infty,\infty)\coloneqq\lim_{k\to\infty}\tau(k,k)$ exists. By Theorem 5.1, we have $\tau(\infty,\infty)=\tau^{\star}$ if Assumptions 3.1 and 3.2 hold.
Provided the presence of the correlative sparsity pattern stated in Assumption 3.3, sparse SDP relaxations for such a problem can be constructed as follows:
$\left\{\begin{aligned} \tau^{\mbox{\tiny cs}}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})\coloneqq\sup_{\gamma,\sigma_{0},S_{j}}&\ \gamma\\ \text{s.t.}&\ f({\boldsymbol{x}})-\gamma=\sum_{\ell=1}^{t}\left(\sigma_{\ell,0}({\boldsymbol{x}}(\mathcal{I}_{\ell}))+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\left\langle S_{j}({\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}),G_{j}({\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})\right),\\ &\ \sigma_{\ell,0}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})],S_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}]^{q_{j}},\ j\in\mathcal{J}_{\ell},\ \ell\in[t],\,\text{are SOS},\\ &\ \deg\sigma_{\ell,0}\leq 2k_{{\boldsymbol{x}}},\ \deg_{{\boldsymbol{x}}}S_{j}\leq 2(k_{{\boldsymbol{x}}}-\lceil\deg_{{\boldsymbol{x}}}G_{j}/2\rceil),\ \deg_{{\boldsymbol{y}}}S_{j}\leq 2(k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G_{j}/2\rceil).\end{aligned}\right$. (12)
For the sequence $\{\tau^{\mbox{\tiny cs}}(k,k)\}_{k\in\mathbb{N}}$, if the assumptions in Theorem 3.2 hold, we have the convergence
$\tau^{\mbox{\tiny cs}}(\infty,\infty)\coloneqq\lim_{k\to\infty}\tau^{\mbox{\tiny cs}}(k,k)=\tau^{\star}.$
When the number $\max\{|\mathcal{I}_{\ell}|,\ell\in[t]\}$ is small, dramatic computational savings can be expected.
Example 5.2.
Consider the problem $f^{\star}\coloneqq\min_{{\boldsymbol{x}}\in\widehat{X}}f({\boldsymbol{x}})$ where $f({\boldsymbol{x}})=\sum_{i=1}^{n}(x_{i}+1)^{4}$,
$\widehat{\mathcal{X}}=\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G_{j}({\boldsymbol{x}},y)\succeq 0,\ \forall j\in[n-1],\ y\in\mathcal{Y}\},\quad G_{j}({\boldsymbol{x}},y)=\left[\begin{array}[]{cc}1-yx_{j}&x_{j+1}\\ x_{j+1}&1+yx_{j}\end{array}\right],\ j\in[n-1],$
and $\mathcal{Y}=[0,1]$. Clearly, Assumption 3.3 is satisfied with t=n1t=n-1, $\mathcal{I}_{\ell}=\{\ell,\ell+1\}$ and $\mathcal{J}_{\ell}=\{\ell\}$ for $\ell\in[n-1]$. It is easy to see that
$\widehat{\mathcal{X}}=\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid x_{i}^{2}+x_{i+1}^{2}\leq 1,\ i\in[n-1]\}.$
Since the problem $\min_{{\boldsymbol{x}}\in\widehat{X}}f({\boldsymbol{x}})$ is convex, numerical values for $f^{\star}$ can be obtained by solving the associated KKT system. These results can then be used to evaluate the accuracy of the SDP lower bounds $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tau^{\mbox{\tiny cs}}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$. We choose $\nu$ in the SDP relaxations (11) and (12) to be the Lebesgue measure on $\mathcal{Y}$ and present the numerical results in Table 3, where the symbol ‘–/–’ indicates that MOSEK runs out of memory. From the table, we could see that dramatic computational advantages are gained by exploiting correlative sparsity. ∎
Table 3. Computational results for Example 5.2.
$n$ $f^{\star}$ $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}$)=(2, 10) $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}$)=(2, 15)
$\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tau^{\mbox{\tiny cs}}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tau^{\mbox{\tiny cs}}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time
33 0.0206 0.0184/13.1s 0.0184/4.45s 0.0191/83.3s 0.0191/19.9s
44 0.0294 0.0262/65.3s 0.0262/9.23s –/– 0.0273/36.5s
55 0.0359 –/– 0.0320/13.9s –/– 0.0333/45.1s
66 0.0441 –/– 0.0394/19.2s –/– 0.0409/62.5s
5.3. The non-compact case
Let us consider the problem (8) with a non‑compact set $\mathcal{X}$. For such a problem, after fixing a constant $\varepsilon>0$, we can apply Corollary 4.1 to construct the following hierarchy of SDP relaxations:
$\left\{\begin{aligned} \tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})\coloneqq\inf_{{\boldsymbol{\gamma}},S_{0},S}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ \theta^{k_{{\boldsymbol{x}}}-d_{\Lambda}}(\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})+\varepsilon\theta^{d_{\Lambda}}I_{p})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\ \deg S_{0}\leq 2k_{{\boldsymbol{x}}},\ \deg_{{\boldsymbol{x}}}S\leq 2(k_{{\boldsymbol{x}}}-d_{G}),\ \deg_{{\boldsymbol{y}}}S\leq 2(k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil),\end{aligned}\right$. (13)
where $d_{\Lambda}\coloneqq\max_{i,j\in[p]}\lfloor\deg_{{\boldsymbol{x}}}(\Lambda_{ij})/2\rfloor+1$. Let $\mathcal{F}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ denote the feasible set of (13).
For the sequence $\{\tau_{\varepsilon}(\ell,\ell)\}_{\varepsilon>0,\ell\in\mathbb{N}}$, let $\tau_{0}(\infty,\infty)\coloneqq\lim_{\varepsilon\to 0}\lim_{\ell\to\infty}\tau_{\varepsilon}(\ell,\ell)$.
Theorem 5.2.
Let Assumption 3.2 hold. Then the limit $\tau_{0}(\infty,\infty)$ exists, and $\tau_{0}(\infty,\infty)\leq\tau^{\star}$. Moreover, if the optimal solution set $\mathcal{S}$ of (8) is nonempty and bounded, we have $\tau_{0}(\infty,\infty)=\tau^{\star}$.
Proof.
Since Assumption 3.2 holds, Corollary 4.1 implies that, for any $\varepsilon>0$ and sufficiently large $\ell$, $\mathcal{F}_{\varepsilon}(\ell,\ell)$ contains the feasible set $\mathcal{F}$ of (8) and hence $\tau_{\varepsilon}(\ell,\ell)\leq\tau^{\star}$. It is clear that $\tau_{\varepsilon}(\ell,\ell)$ is non-increasing as $\ell\to\infty$ for fixed $\varepsilon>0$ and is non-decreasing as $\varepsilon\to 0$ for fixed $\ell$. Hence, the limit $\tau_{0}(\infty,\infty)$ exists and $\tau_{0}(\infty,\infty)\in[-\infty,\tau^{\star}]$.
Assuming that the optimal solution set $\mathcal{S}$ of (8) is nonempty and bounded, we show that $\tau_{0}(\infty,\infty)=\tau^{\star}$. Suppose on the contrary that $\tau_{0}(\infty,\infty)<\tau^{\star}$. Then, there exists a sequence $\{\varepsilon_{j}\}_{j\in\mathbb{N}}$ with $\lim_{j\to\infty}\varepsilon_{j}=0$ and a sequence $\{{\boldsymbol{\gamma}}^{(\ell_{j})}\}_{j\in\mathbb{N}}$ of feasible points of (13) with $(\varepsilon,k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})=(\varepsilon_{j},\ell_{j},\ell_{j})$ such that for each $j\in\mathbb{N}$, ${\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}^{(\ell_{j})}\leq\tau^{\star}-\delta$ for some $\delta>0$. We claim that $\{{\boldsymbol{\gamma}}^{(\ell_{j})}\}_{j\in\mathbb{N}}$ is bounded. Suppose that $\|{\boldsymbol{\gamma}}^{(\ell_{j})}\|\to\infty$ as $j\to\infty$ and without loss of generality that $\lim_{j\to\infty}{\boldsymbol{\gamma}}^{(\ell_{j})}/\|{\boldsymbol{\gamma}}^{(\ell_{j})}\|={\boldsymbol{v}}$ for some ${\boldsymbol{v}}=(v_{1},\ldots,v_{r})\in{\mathbb{R}}^{r}$ with $\|{\boldsymbol{v}}\|=1$. Since ${\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}^{(\ell_{j})}$ is bounded by $\tau^{\star}-\delta$ for all $j\in\mathbb{N}$, we have ${\boldsymbol{c}}^{\intercal}{\boldsymbol{v}}=0$. Moreover, for arbitrarily fixed ${\boldsymbol{u}}\in\mathcal{X}$, we have
$\Lambda({\boldsymbol{u}},{\boldsymbol{\gamma}}^{(\ell_{j})})+\varepsilon_{j}\theta({\boldsymbol{u}})^{d_{\Lambda}}I_{p}=\Lambda_{0}({\boldsymbol{u}})-\sum_{i=1}^{r}\Lambda_{i}({\boldsymbol{u}})\gamma^{(\ell_{j})}_{i}+\varepsilon_{j}\theta({\boldsymbol{u}})^{d_{\Lambda}}I_{p}\succeq 0,\ \forall j\in\mathbb{N}.$ (14)
Dividing the left-hand side of (14) by $\|{\boldsymbol{\gamma}}^{(\ell_{j})}\|$ and letting $j\to\infty$, we get $\sum_{i=1}^{r}\Lambda_{i}({\boldsymbol{u}})v_{i}\succeq 0$. Since $\boldsymbol{u}$} is an arbitrary point in $\mathcal{X}$ and ${\boldsymbol{c}}^{\intercal}{\boldsymbol{v}}=0$, for any point ${\boldsymbol{\gamma}}^{\star}\in\mathcal{S}$, we have ${\boldsymbol{\gamma}}^{\star}+t{\boldsymbol{v}}\in\mathcal{S}$ for all $t>0$, which contradicts the boundedness of $\mathcal{S}$. As we have proved that $\{{\boldsymbol{\gamma}}^{(\ell_{j})}\}_{j\in\mathbb{N}}$ is bounded, we can assume without loss of generality that $\lim_{j\to\infty}{\boldsymbol{\gamma}}^{(\ell_{j})}=\hat{{\boldsymbol{\gamma}}}$ for some $\hat{{\boldsymbol{\gamma}}}\in{\mathbb{R}}^{r}$. Then letting $j\to\infty$ in (14) yields the result that $\hat{{\boldsymbol{\gamma}}}$ is feasible to (8). Thus, we have
$\tau^{\star}\leq{\boldsymbol{c}}^{\intercal}\hat{{\boldsymbol{\gamma}}}=\lim_{j\to\infty}{\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}^{(\ell_{j})}\leq\tau^{\star}-\delta,$
a contradiction. ∎
Now, let us consider the problem
$\tilde{\tau}^{\star}\coloneqq\inf_{{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}}\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\ \text{ s.t. }\ \Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})\succeq 0,\ \forall{\boldsymbol{x}}\in\mathcal{X}\cap{\mathbb{R}}_{+}^{n},$ (15)
where $\mathcal{X}$ is non-compact and defined in (8), and $\mathcal{Y}$ is defined in (7). After fixing a constant $\varepsilon>0$, we apply Corollary 4.3 to construct the following hierarchy of SDP relaxations:
$\left\{\begin{aligned} \tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})\coloneqq\inf_{{\boldsymbol{\gamma}},S_{0},S}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ (1+{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{2k_{{\boldsymbol{x}}}-2d_{\Lambda}}(\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})+\varepsilon(1+{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{2d_{\Lambda}}I_{p})=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ P_{0}=\sum_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2k_{{\boldsymbol{x}}}}}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}\text{ with each }P_{0,{\boldsymbol{\alpha}}}\succeq 0,\\ &\ P=\sum_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2(k_{{\boldsymbol{x}}}-d_{G})}}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq},\ \deg_{{\boldsymbol{y}}}P\leq 2(k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil),\\ &\ \text{ with each }P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\in\mathcal{Q}^{pq}_{k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil}(\mathbf{H}).\end{aligned}\right$. (16)
For the sequence $\{\tilde{\tau}_{\varepsilon}(\ell,\ell)\}_{\varepsilon>0,\ell\in\mathbb{N}}$, let $\tilde{\tau}_{0}(\infty,\infty)\coloneqq\lim_{\varepsilon\to 0}\lim_{\ell\to\infty}\tilde{\tau}_{\varepsilon}(\ell,\ell)$. Recall the set $\mathbf{H}$ in Corollary 4.2. Similarly to Theorem 5.2, we have the following convergence result.
Theorem 5.3.
Let the Archimedean condition hold for the quadratic module $\mathcal{Q}^{pq}(\mathbf{H})$. Then the limit $\tilde{\tau}_{0}(\infty,\infty)$ exists and $\tilde{\tau}_{0}(\infty,\infty)\leq\tau^{\star}$. Moreover, if the optimal solution set $\mathcal{S}$ of (15) is nonempty and bounded, we have $\tilde{\tau}_{0}(\infty,\infty)=\tau^{\star}$.
Note that we set the same degree bounds on $\boldsymbol{x}$} and $\boldsymbol{y}$} in the representations of $\Lambda({\boldsymbol{x}},{\boldsymbol{y}})$ in (13) and (16), so that we can compare the behaviors of $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ when applying both (13) and (16) to (8) if $\mathcal{X}=\mathcal{X}\cap{\mathbb{R}}^{n}_{+}$.
Example 5.3.
Consider the problem (8) where we let $m=1$, $n=r=2$, $\mathcal{Y}=[0,1]$,
$G({\boldsymbol{x}},y)=Q_{1}^{\intercal}\left[\begin{array}[]{cc}x_{1}+x_{2}&0\\ 0&-2y(1-y)+(1-y)^{2}x_{1}+y^{2}x_{2}\end{array}\right]Q_{1},$
and
$\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})=Q_{2}^{\intercal}\left[\begin{array}[]{cc}-2+\left(-\frac{\gamma_{1}}{2}+\frac{3\gamma_{2}}{2}+3\right)x_{2}+\left(\frac{3\gamma_{1}}{2}-\frac{\gamma_{2}}{2}+3\right)x_{1}&0\\ 0&-2-\gamma_{1}x_{2}-\gamma_{2}x_{1}\end{array}\right]Q_{2}.$
Here, $Q_{1}$ and $Q_{2}$ are two fixed arbitrary $2\times 2$ orthogonal matrices.
For any fixed $(x_{1},x_{2})$, consider the function
$f(y)=-2y(1-y)+(1-y)^{2}x_{1}+y^{2}x_{2}=(x_{1}+x_{2}+2)y^{2}-2(x_{1}+1)y+x_{1}.$
The evaluations $f(0)=x_{1}$ and $f(1)=x_{2}$ imply that $x_{1}\geq 0$ and $x_{2}\geq 0$. Assuming $x_{1},x_{2}\geq 0$, the minimum of $f(y)$ is
$f\left(\frac{x_{1}+1}{x_{1}+x_{2}+2}\right)=\frac{x_{1}x_{2}-1}{x_{1}+x_{2}+2}\ \text{ where }\ \frac{x_{1}+1}{x_{1}+x_{2}+2}\in\mathcal{Y}.$
Thus we have
$\mathcal{X}=\{(x_{1},x_{2})\in{\mathbb{R}}^{2}\mid x_{1}\geq 0,\ x_{2}\geq 0,\ x_{1}x_{2}\geq 1\}.$
Then similarly, we can show that $\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})\succeq 0$ for all ${\boldsymbol{x}}\in\mathcal{X}\cap{\mathbb{R}}_{+}^{n}$ if and only if
$2+\left(-\frac{\gamma_{1}}{2}+\frac{3\gamma_{2}}{2}+3\right)x_{2}+\left(\frac{3\gamma_{1}}{2}-\frac{\gamma_{2}}{2}+3\right)x_{1}\geq 0\ \text{ and }\ -2-\gamma_{1}x_{2}-\gamma_{2}x_{1}\geq 0$
on $\mathcal{X}$ if and only if
$\frac{\gamma_{1}}{2}-\frac{3\gamma_{2}}{2}-3\leq 0,\ -\frac{3\gamma_{1}}{2}+\frac{\gamma_{2}}{2}-3\leq 0,\ \phi_{1}({\boldsymbol{\gamma}})\coloneqq\left(\frac{\gamma_{1}}{2}-\frac{3\gamma_{2}}{2}-3\right)\left(-\frac{3\gamma_{1}}{2}+\frac{\gamma_{2}}{2}-3\right)-1\geq 0,$
and
$\gamma_{1}\leq 0,\ \gamma_{2}\leq 0,\ \phi_{2}({\boldsymbol{\gamma}})\coloneqq\gamma_{1}\gamma_{2}-1\geq 0.$
The feasible set $\mathcal{F}$ is shown in light blue in Figure 2. Since $\mathcal{X}=\mathcal{X}\cap{\mathbb{R}}^{2}_{+}$, we can apply both SDP relaxations (13) and (16) to this problem and compare the behaviors of $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$.
Consider the objective function ${\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}=\gamma_{1}-\gamma_{2}$. One can check that the optimal solution is one of the solutions of $\phi_{1}({\boldsymbol{\gamma}})=\phi_{2}({\boldsymbol{\gamma}})=0$:
${\boldsymbol{\gamma}}^{\star}=\left(1-\sqrt{5}-\sqrt{5-2\sqrt{5}},1-\sqrt{5}+\sqrt{5-2\sqrt{5}}\right)\approx(-1.9626,-0.5095),$
and the optimal value $\tau^{\star}\approx-1.4531$. The computational results for $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ with different $\varepsilon$ and $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ are presented in Table 4. As expected, for a given $\varepsilon>0$, both $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ provide lower bounds for the optimal value $\tau^{\star}$ when $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$ are sufficiently large (except for $\tau_{0.01}(2,15)$ and $\tilde{\tau}_{0.01}(2,15)$ which requires larger $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$ as $\varepsilon=0.01$ is small). Moreover, for a fixed $\varepsilon>0$, both quantities are non-increasing as $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$ increase, and for fixed $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$, they are non-decreasing as $\varepsilon$ decreases. Compared with $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$, the values $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ with the same $(\varepsilon,k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ are smaller and can be computed in much less time.
To illustrate the behavior of optimal solutions of (13) and (16) under varying $\varepsilon$ and $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$, like in Example 5.1, we solve (13) and (16) with the objective function $\gamma_{1}\cos\theta+\gamma_{2}\sin\theta$ for $40$ evenly spaced values of $\theta$ over $0,2\pi]$, and extract the corresponding optimal solutions. The feasible sets of (13) and (16), which serve as approximations of $\mathcal{F}$, are then outlined by connecting these optimal solutions and shown in Figure 2. As one can see, the behavior of these optimal points aligns with the quantities $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$, confirming the theoretical expectations discussed above. In particular, it is evident from the figure that the feasible set of (16) contains that of (13) with the same $(\varepsilon,k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$, which corresponds to the fact that $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ is smaller than $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$. ∎
Table 4. Computational results for Example 5.3.
$\varepsilon$ $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}$)=(2, 5) $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}$)=(2, 10) $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}$)=(2, 15)
$\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tilde{\tau}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tilde{\tau}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tilde{\tau}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time
$0.1$ -1.3716/2.20s -1.4922/1.69s -1.4844/32.3s -1.6587/7.32s -1.5650/547s -1.6837/27.2s
$0.05$ -1.2664/2.34s -1.3163/1.93s -1.3698/31.1s -1.4899/7.21s -1.4552/507s -1.5230/29.9s
$0.01$ -1.1873/2.57s -1.1806/2.17s -1.2949/33.3s -1.3625/7.86s -1.3721/451s -1.3900/35.7s
Figure 2. The feasible set (in light blue) of the problem in Example 5.3 and the feasible sets of its SDP relaxations (13) (outlined with blue line) and (16) (outlined with red line) with the same values $(\varepsilon,k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$.
6. Proofs
6.1. Proof of Proposition 3.4
For a matrix $N\in{\mathbb{S}}^{q}$, denote by $\lambda_{\max}(N)$ and $\lambda_{\min}(N)$ the largest and smallest eigenvalues of $N$, respectively.
Proof of Proposition 3.4.
For each $j\in[s]$, let
$M\coloneqq\max\left\{|\lambda_{\max}(G_{j}({\boldsymbol{x}},{\boldsymbol{y}}))|,|\lambda_{\min}(G_{j}({\boldsymbol{x}},{\boldsymbol{y}}))|\colon{\boldsymbol{x}}\in[-C,C]^{n},\ {\boldsymbol{y}}\in\mathcal{Y},\ j\in[s]\right\}.$
As $\mathcal{Y}$ is compact, the quantity $M$ is well-defined and positive. For this $M$, we show that there exists some $\bar{k}\in\mathbb{N}$ such that (5) holds for all $k\geq\bar{k}$. Suppose on the contrary that for any $k\in\mathbb{N}$, there exists ${\boldsymbol{x}}^{(k)}\in[-C,C]^{n}$ such that
$F({\boldsymbol{x}}^{(k)})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}}^{(k)},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}}^{(k)},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\not\succ 0.$ (17)
As $C,C]^{n}$ is compact, without loss of generality, we may assume that $\lim_{k\to\infty}{\boldsymbol{x}}^{(k)}={\boldsymbol{x}}^{\star}$ for some ${\boldsymbol{x}}^{\star}\in[-C,C]^{n}$. Next, we claim that there exists $k^{\prime}\in\mathbb{N}$ and a neighborhood $\mathcal{O}_{1}$ of $\boldsymbol{x}$$^{\star}$ such that (5) holds for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$, which yields a contradiction.
We first consider the case that $^{\boldsymbol{x}}^{\star}\in\widehat{\mathcal{X}}$. As $F({\boldsymbol{x}})\succ 0$ on $\widehat{\mathcal{X}}$, there exists $\varepsilon>0$ and a neighborhood $\mathcal{O}_{1}$ of $\boldsymbol{x}$$^{\star}$ such that $F({\boldsymbol{x}})\succeq\varepsilon I_{p}$ for all ${\boldsymbol{x}}\in\mathcal{O}_{1}$. We now prove that there exsits $k^{\prime}\in\mathbb{N}$ such that for all $k\geq k^{\prime}$, it holds
$\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})\leq\frac{\varepsilon}{2},$
for all ${\boldsymbol{x}}\in[-C,C]^{n}$. Fix a pair of ${\boldsymbol{u}}\in[-C,C]^{n}$ and ${\boldsymbol{w}}\in\mathcal{Y}$. Let $\{\lambda^{(j)}_{i}\}_{i\in[q_{j}]}$ be the set of eigenvalues of $G_{j}({\boldsymbol{u}},{\boldsymbol{w}})$. Take the decomposition $G_{j}({\boldsymbol{u}},{\boldsymbol{w}})=Q_{j}D_{j}Q_{j}^{\intercal}$ where $D_{j}=\mbox{diag}(\lambda_{1}^{(j)},\ldots,\lambda^{(j)}_{q_{j}})$ and $Q_{j}\in{\mathbb{R}}^{q_{j}\times q_{j}}$ with $Q_{j}^{\intercal}Q_{j}=Q_{j}Q_{j}^{\intercal}=I_{q_{j}}$. Then,
$\begin{aligned} \left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{u}},{\boldsymbol{w}})/M\right)^{2k},G_{j}({\boldsymbol{u}},{\boldsymbol{w}})\right\rangle=&\left\langle\left(I_{q_{j}}-Q_{j}(D_{j}/M)Q_{j}^{\intercal}\right)^{2k},Q_{j}D_{j}Q_{j}^{\intercal}\right\rangle\\ =&\left\langle Q_{j}\left(I_{q_{j}}-D_{j}/M\right)^{2k}Q_{j}^{\intercal},Q_{j}D_{j}Q_{j}^{\intercal}\right\rangle=M\sum_{i=1}^{q_{j}}\frac{\lambda_{i}^{(j)}}{M}\left(1-\frac{\lambda_{i}^{(j)}}{M}\right)^{2k}\end{aligned}.$ (18)
Let $q^{\prime}\coloneqq\max_{j\in[s]}q_{j}$. Note that there exists $k^{\prime}\in\mathbb{N}$ such that for all $k\geq k^{\prime}$,
$\max\left\{\xi(1-\xi)^{2k}:\xi\in[0,1]\right\}\leq\frac{\varepsilon}{2q^{\prime}sM\nu(\mathcal{Y})}.$
Hence, for any pair of ${\boldsymbol{x}}\in[-C,C]^{n}$ and ${\boldsymbol{y}}\in\mathcal{Y}$, from (18) we obtain that
$\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\leq Mq_{j}\frac{\varepsilon}{2q^{\prime}sM\nu(\mathcal{Y})}\leq\frac{\varepsilon}{2s\nu(\mathcal{Y})}$ (19)
for all $k\geq k^{\prime}$, where $k^{\prime}$ does not depends on the choice of ${\boldsymbol{x}}\in[-C,C]^{n}$ and ${\boldsymbol{y}}\in\mathcal{Y}$. Therefore, for all ${\boldsymbol{x}}\in[-C,C]^{n}$ and $k\geq k^{\prime}$, we have
$\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})\leq\sum_{j=1}^{s}\int_{\mathcal{Y}}\frac{\varepsilon}{2s\nu(\mathcal{Y})}\mathrm{d}\nu({\boldsymbol{y}})\leq\frac{\varepsilon}{2}.$
Then, for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$,
$F({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\succeq\varepsilon I_{p}-\frac{\varepsilon}{2}I_{p}=\frac{\varepsilon}{2}I_{p}\succ 0.$
Next, we consider the case that $^{\boldsymbol{x}}^{\star}\in[-C,C]^{n}\setminus\widehat{\mathcal{X}}$. There exists $j_{0}\in[s]$ and ${\boldsymbol{y}}^{(0)}\in\mathcal{Y}$ such that $\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}}^{\star},{\boldsymbol{y}}^{(0)})<0$. By continuity, for some $\bar{\lambda}<0$, there exists a neighborhood $\mathcal{O}_{1}$ (resp., $\mathcal{O}_{2}$) of $\boldsymbol{x}$$^{\star}$ (resp., ${\boldsymbol{y}}^{(0)}$) such that $\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}}))\leq\bar{\lambda}$ for all ${\boldsymbol{x}}\in\mathcal{O}_{1}$ and ${\boldsymbol{y}}\in\mathcal{O}_{2}$. For any pair of ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and ${\boldsymbol{y}}\in\mathcal{O}_{2}\cap\mathcal{Y}$, letting $\lambda_{\min}^{(j_{0})}\coloneqq\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}}))$, from (18) we obtain that
$\left\langle\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle$ $\leq\sum_{\lambda_{i}^{(j_{0})}>0}\lambda_{i}^{(j_{0})}+\sum_{\lambda_{i}^{(j_{0})}<0}\lambda_{i}^{(j_{0})}\left(1-\frac{\lambda_{i}^{(j_{0})}}{M}\right)^{2k}$ (20)
$\leq Mq^{\prime}+\lambda_{\min}^{(j_{0})}\left(1-\frac{\lambda_{\min}^{(j_{0})}}{M}\right)^{2k}\leq Mq^{\prime}+\bar{\lambda}\left(1-\frac{\bar{\lambda}}{M}\right)^{2k}.$
Since $\mathrm{supp}(\nu)=\mathcal{Y}$, we have $\nu(\mathcal{Y}\cap\mathcal{O}_{2})>0$. Note that for each $j\in[s]$, by (19), the maximal eigenvalue of
$\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}=\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle I_{p}$
could be uniformly bounded from above on $C,C]^{n}\times\mathcal{Y}$ for all $k\in\mathbb{N}$. Therefore, as $C,C]^{n}$ and $\mathcal{Y}$ are compact, there exists $R\in{\mathbb{R}}$ such that
$F({\boldsymbol{x}})-\sum_{j\neq j_{0}}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$ (21)
$-\int_{\mathcal{Y}\setminus\mathcal{O}_{2}}\left\langle I_{p}\otimes\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})-(Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1)I_{p}\succeq RI_{p}$
for all ${\boldsymbol{x}}\in[-C,C]^{n}$. Since $1-\bar{\lambda}/M>1$, there exists $k^{\prime}\in\mathbb{N}$ such that
$\bar{\lambda}\left(1-\frac{\bar{\lambda}}{M}\right)^{2k}<\frac{R}{\nu(\mathcal{Y}\cap\mathcal{O}_{2})}$ (22)
holds for all $k\geq k^{\prime}$. Then, for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$,
$F({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
$\succeq$ $-\int_{\mathcal{Y}\cap\mathcal{O}_{2}}\left\langle I_{p}\otimes\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})+(Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R)I_{p}$ (by (21))
$\succeq$ $\left(-\int_{\mathcal{Y}\cap\mathcal{O}_{2}}\left(Mq^{\prime}+\bar{\lambda}\left(1-\frac{\bar{\lambda}}{M}\right)^{2k}\right)\mathrm{d}\nu({\boldsymbol{y}})+Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R\right)I_{p}$ (by (20))
$\succeq$ $\left(-\left(Mq^{\prime}+\frac{R}{\nu(\mathcal{Y}\cap\mathcal{O}_{2})}\right)\nu(\mathcal{Y}\cap\mathcal{O}_{2})+Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R\right)I_{p}=I_{p},$ (by (22))
which validates the claim and completes the proof. ∎
Using Proposition 3.4, we can recover Theorem 3.1 in the case that $\mathcal{Y}$ is compact.
Lemma 6.1.
For any $h\in\mathcal{Q}^{1}(\mathbf{G},\nu)$ and SOS matrix $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, we have $S({\boldsymbol{x}})h({\boldsymbol{x}})\in\mathcal{Q}^{p}(\mathbf{G},\nu)$.
Proof.
Write $h=\sigma_{0}+\sum_{j=1}^{s}\int_{\mathcal{Y}}\langle S_{j},G_{j}\rangle\mathrm{d}\nu({\boldsymbol{y}})$ where all $\sigma_{0}\in{\mathbb{R}}[{\boldsymbol{x}}]$ and $S_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q_{j}}$ are SOS. Then
$Sh=S\sigma_{0}+\sum_{j=1}^{s}S\int_{\mathcal{Y}}\langle S_{j},G_{j}\rangle\mathrm{d}\nu({\boldsymbol{y}})=S\sigma_{0}+\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle S\otimes S_{j},G_{j}\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
As $S\sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $S\otimes S_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq_{j}}$ are all SOS matrices, we have $Sh\in\mathcal{Q}^{p}(\mathbf{G},\nu)$. ∎
Corollary 6.1.
Suppose that $\mathcal{Y}$ is compact and Assumption 3.1 holds. If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\widehat{\mathcal{X}}$, then $F\in\mathcal{Q}^{p}(\mathbf{G},\nu)$.
Proof.
Let $C>0$ be given in Assumption 3.1. Then, by Proposition 3.4, there exists $M>0$ and $\bar{k}\in\mathbb{N}$ such that
$F({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\succ 0$
holds on the set $\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid\|{\boldsymbol{x}}\|^{2}\leq C\}$ for all $k\geq\bar{k}$. By Scherer-Hol’s Positivestellensatz (Theorem 2.1), there exist SOS matrices $S_{0},S_{1}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ such that
$F({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})=S_{0}({\boldsymbol{x}})+S_{1}({\boldsymbol{x}})(C-\|{\boldsymbol{x}}\|^{2}).$
By Lemma 6.1, $S_{1}({\boldsymbol{x}})(C-\|{\boldsymbol{x}}\|^{2})\in\mathcal{Q}^{p}(\mathbf{G},\nu)$ which implies that $F\in\mathcal{Q}^{p}(\mathbf{G},\nu)$. ∎
6.2. Proof of Theorem 4.1
Proof of Theorem 4.1.
Let
$\widetilde{\mathcal{X}}\coloneqq\left\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid 1-\|{\boldsymbol{x}}\|^{2}=0,\ G({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\right\}.$
Then $F({\boldsymbol{u}})\succ 0$ for all ${\boldsymbol{u}}\in\widetilde{\mathcal{X}}$. By Corollary 3.1, it holds that
$F({\boldsymbol{x}})=S^{\prime}_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S^{\prime}({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})+H({\boldsymbol{x}})(1-\|{\boldsymbol{x}}\|^{2}),$
where $^{\prime}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$^{\prime}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ are SOS matrices and $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$. Replacing $\boldsymbol{x}$} by ${\boldsymbol{x}}/\|{\boldsymbol{x}}\|$ in the above equality yields
$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{-\deg{F}}=S^{\prime}_{0}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|}\right)+\int_{\mathcal{Y}}\left\langle S^{\prime}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right),G\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Let
$k^{\prime}\coloneqq\max\left\{\deg{F},\ \deg S^{\prime}_{0},\ \deg_{{\boldsymbol{x}}}{G}+\deg_{{\boldsymbol{x}}}S^{\prime}\right\}.$
By assumption, $k^{\prime}$ is even. Multiplying the two sides of the last equality with $\|{\boldsymbol{x}}\|^{k^{\prime}}$ gives
$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{k^{\prime}-\deg{F}}=\overline{S}_{0}\left({\boldsymbol{x}}\right)+\int_{\mathcal{Y}}\left\langle\overline{S}\left({\boldsymbol{x}},{\boldsymbol{y}}\right),G\left({\boldsymbol{x}},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),$ (23)
where
$\overline{S}_{0}\coloneqq S^{\prime}_{0}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|}\right)\|{\boldsymbol{x}}\|^{k^{\prime}}\ \ \text{and}\ \ \overline{S}\coloneqq S^{\prime}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right)\|{\boldsymbol{x}}\|^{k^{\prime}-\deg_{{\boldsymbol{x}}}{G}}.$
Since $S^{\prime}$ is an SOS matrix and $k^{\prime}-\deg_{{\boldsymbol{x}}}{G}\geq\deg_{{\boldsymbol{x}}}S^{\prime}$, we have $\overline{S}=H^{\intercal}H$ with $H=H_{1}+H_{2}\|{\boldsymbol{x}}\|$ where $H_{1},H_{2}\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{\ell\times pq}$ for some $\ell\in\mathbb{N}$, are homogeneous in $\boldsymbol{x}$} of degree $(k^{\prime}-\deg_{{\boldsymbol{x}}}{G})/2$ and $(k^{\prime}-\deg_{{\boldsymbol{x}}}{G})/2-1$, respectively. Thus,
$\overline{S}=H^{\intercal}H=(H_{1}+H_{2}\|{\boldsymbol{x}}\|)^{\intercal}(H_{1}+H_{2}\|{\boldsymbol{x}}\|)=\left(H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}\right)+\left(H_{1}^{\intercal}H_{2}+H_{2}^{\intercal}H_{1}\right)\|{\boldsymbol{x}}\|.$
Similarly, there exist homogeneous polynomial matrices $H_{0,1}\in{\mathbb{R}}[x]^{\ell_{0}\times p}_{k^{\prime}/2}$ and $H_{0,2}\in{\mathbb{R}}[{\boldsymbol{x}}]^{\ell_{0}\times p}_{k^{\prime}/2-1}$ for some $\ell_{0}\in\mathbb{N}$ such that
$\overline{S}_{0}=\left(H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}\right)+\left(H_{0,1}^{\intercal}H_{0,2}+H_{0,2}^{\intercal}H_{0,1}\right)\|{\boldsymbol{x}}\|.$
Then, by (23), it holds that
$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{k^{\prime}-\deg{F}}=$ $\left(H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}\right)+\int_{\mathcal{Y}}\left\langle\left(H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}\right),G\left({\boldsymbol{x}},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
$+\left(\left(H_{0,1}^{\intercal}H_{0,2}+H_{0,2}^{\intercal}H_{0,1}\right)+\int_{\mathcal{Y}}\left\langle\left(H_{1}^{\intercal}H_{2}+H_{2}^{\intercal}H_{1}\right),G\left({\boldsymbol{x}},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\|{\boldsymbol{x}}\|.$
Since $\|{\boldsymbol{x}}\|$ is not a polynomial and the left hand side of the above equation is a polynomial matrix, we must have
$\left(H_{0,1}^{\intercal}H_{0,2}+H_{0,2}^{\intercal}H_{0,1}\right)+\int_{\mathcal{Y}}\left\langle\left(H_{1}^{\intercal}H_{2}+H_{2}^{\intercal}H_{1}\right),G\left({\boldsymbol{x}},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})=0.$
Then, letting $N\coloneqq(k^{\prime}-\deg{F})/2\in\mathbb{N}$, $S_{0}\coloneqq H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}$ and $S\coloneqq H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}$, we obtain the desired conclusion. ∎
6.3. Proof of Theorem 4.2
Let ${\boldsymbol{z}}=(z_{1},\ldots,z_{n})$ and ${\boldsymbol{z}}\circ{\boldsymbol{z}}=(z_{1}^{2},\ldots,z_{n}^{2})$. We first derive an intermediary result.
Theorem 6.1.
Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$}, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $N\in\mathbb{N}$ and polynomial matrices $_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in $\boldsymbol{x}$} ($\deg S_{0}=N+\deg F$ and $\deg_{{\boldsymbol{x}}}S=N+\deg F-\deg_{{\boldsymbol{x}}}G$), and satisfy that $S_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$ (resp., $S({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$) are SOS matrices in $\boldsymbol{z}$} (resp., $\boldsymbol{z}$} and $\boldsymbol{y}$}), such that
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Proof.
Consider the polynomial matrices $F({\boldsymbol{z}}\circ{\boldsymbol{z}})\in{\mathbb{S}}[{\boldsymbol{z}}]^{p}$ and $G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{z}},{\boldsymbol{y}}]^{q}$. Then, by assumption and Theorem 4.1, there exists $N$ and polynomial matrices $H_{0}\in{\mathbb{R}}[{\boldsymbol{z}}]^{\ell_{0}\times p}$, $H\in{\mathbb{R}}[{\boldsymbol{z}},{\boldsymbol{y}}]^{\ell\times pq}$ for some $\ell_{0},\ell\in\mathbb{N}$ that are homogeneous in $\boldsymbol{z}$}, such that
$\|{\boldsymbol{z}}\|^{2N}F({\boldsymbol{z}}\circ{\boldsymbol{z}})=(H_{0}^{\intercal}H_{0})({\boldsymbol{z}})+\int_{\mathcal{Y}}\left\langle(H^{\intercal}H)({\boldsymbol{z}},{\boldsymbol{y}}),G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Moreover, by Theorem 4.1, $\deg H_{0}=N+\deg F$, and $\deg_{{\boldsymbol{x}}}H=N+\deg F-\deg_{{\boldsymbol{x}}}G$. Note that there are sets of polynomial matrices
$\{H_{0,{\boldsymbol{\alpha}}}({\boldsymbol{x}}):{\boldsymbol{\alpha}}\in\{0,1\}^{n}\}\subset{\mathbb{R}}[{\boldsymbol{x}}]^{\ell_{0}\times p}\quad\text{and}\quad\{H_{{\boldsymbol{\alpha}}}({\boldsymbol{x}},{\boldsymbol{y}}):{\boldsymbol{\alpha}}\in\{0,1\}^{n}\}\subset{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{\ell\times p}$
which are homogeneous in $\boldsymbol{x}$}, such that
$H_{0}({\boldsymbol{z}})=\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{z}}^{{\boldsymbol{\alpha}}}H_{0,{\boldsymbol{\alpha}}}({\boldsymbol{z}}\circ{\boldsymbol{z}}),\quad H({\boldsymbol{z}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{z}}^{{\boldsymbol{\alpha}}}H_{{\boldsymbol{\alpha}}}({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}}).$
Then
$({\boldsymbol{e}}^{\intercal}({\boldsymbol{z}}\circ{\boldsymbol{z}}))^{N}F({\boldsymbol{z}}\circ{\boldsymbol{z}})$
$=$ $\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}({\boldsymbol{z}}\circ{\boldsymbol{z}})^{{\boldsymbol{\alpha}}}(H_{0,{\boldsymbol{\alpha}}}^{\intercal}H_{0,{\boldsymbol{\alpha}}})({\boldsymbol{z}}\circ{\boldsymbol{z}})+\int_{\mathcal{Y}}\left\langle\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}({\boldsymbol{z}}\circ{\boldsymbol{z}})^{{\boldsymbol{\alpha}}}(H_{{\boldsymbol{\alpha}}}^{\intercal}H_{{\boldsymbol{\alpha}}})({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}}),G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
$+$ $\sum_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\{0,1\}^{n},{\boldsymbol{\alpha}}\neq{\boldsymbol{\beta}}}{\boldsymbol{z}}^{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}(H_{0,{\boldsymbol{\alpha}}}^{\intercal}H_{0,{\boldsymbol{\beta}}})({\boldsymbol{z}}\circ{\boldsymbol{z}})+\int_{\mathcal{Y}}\left\langle\sum_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\{0,1\}^{n},{\boldsymbol{\alpha}}\neq{\boldsymbol{\beta}}}{\boldsymbol{z}}^{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}(H_{{\boldsymbol{\alpha}}}^{\intercal}H_{{\boldsymbol{\beta}}})({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}}),G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Comparing the even and odd terms in $\boldsymbol{z}$} in the above equation, we get
$({\boldsymbol{e}}^{\intercal}({\boldsymbol{z}}\circ{\boldsymbol{z}}))^{N}F({\boldsymbol{z}}\circ{\boldsymbol{z}})$
$=$ $\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}({\boldsymbol{z}}\circ{\boldsymbol{z}})^{{\boldsymbol{\alpha}}}(H_{0,{\boldsymbol{\alpha}}}^{\intercal}H_{0,{\boldsymbol{\alpha}}})({\boldsymbol{z}}\circ{\boldsymbol{z}})+\int_{\mathcal{Y}}\left\langle\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}({\boldsymbol{z}}\circ{\boldsymbol{z}})^{{\boldsymbol{\alpha}}}(H_{{\boldsymbol{\alpha}}}^{\intercal}H_{{\boldsymbol{\alpha}}})({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}}),G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
It follows that the equality
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}H_{0,{\boldsymbol{\alpha}}}^{\intercal}({\boldsymbol{x}})H_{0,{\boldsymbol{\alpha}}}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}H_{{\boldsymbol{\alpha}}}^{\intercal}({\boldsymbol{x}},{\boldsymbol{y}})H_{{\boldsymbol{\alpha}}}({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
holds for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}$. As ${\mathbb{R}}^{n}_{+}$ has interior points, the polynomial matrices on the left and right side of the above equality are identical. Letting
$S_{0}({\boldsymbol{x}})\coloneqq\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}H_{0,{\boldsymbol{\alpha}}}^{\intercal}({\boldsymbol{x}})H_{0,{\boldsymbol{\alpha}}}({\boldsymbol{x}})\text{ and }S({\boldsymbol{x}},{\boldsymbol{y}})\coloneqq\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}H_{{\boldsymbol{\alpha}}}^{\intercal}({\boldsymbol{x}},{\boldsymbol{y}})H_{{\boldsymbol{\alpha}}}({\boldsymbol{x}},{\boldsymbol{y}}),$
the conclusion follows. ∎
As an extension of Pólya’s result [35], Scherer and Hol [42] provided the following certificate for homogeneous polynomial matrices being positive on the nonnegative orthant.
Theorem 6.2.
[42, Theorem 3] Suppose that the polynomial matrix $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is homogeneous and $H({\boldsymbol{x}})\succeq\lambda I_{p}$ for some $\lambda>0$ on $\{{\boldsymbol{x}}\in{\mathbb{R}}_{+}^{n}\mid{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}}=1\}$. Write $H({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(H)}H_{{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ with each $H_{{\boldsymbol{\alpha}}}\in{\mathbb{S}}^{p}$ and let
$L(H)\coloneqq\max_{{\boldsymbol{\alpha}}\in\mathrm{supp}(H)}\frac{{\boldsymbol{\alpha}}!}{\deg H!}\|H_{{\boldsymbol{\alpha}}}\|,$
where $\|\cdot\|$ denotes the spectral norm. Then, for all
$N\geq\frac{\deg H(\deg(H)-1)L(H)}{2\lambda}-\deg H,$
the coefficient of ${\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ in $({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}H({\boldsymbol{x}})$ for each ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{N+\deg H}$ is PD.
We will use this theorem as another intermediary result. Let $D\coloneqq\max\left\{\deg F,\deg_{{\boldsymbol{x}}}G\right\}$.
Lemma 6.2.
Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$} and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $\varepsilon>0$ such that the homogeneous polynomial matrix
$F_{\varepsilon}({\boldsymbol{x}})\coloneqq({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg F}F({\boldsymbol{x}})-\varepsilon\left(({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D}I_{p}+\int_{\mathcal{Y}}\left\langle({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg_{{\boldsymbol{x}}}G}I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\succ 0$
for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$.
Proof.
By homogeneity, we only need to prove $F_{\varepsilon}\succ 0$ on $\overline{\mathcal{X}}\coloneqq{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}}=1\}.$ That is, there exists $\varepsilon>0$ such that $F({\boldsymbol{x}})-\varepsilon\left(I_{p}+\int_{\mathcal{Y}}\left\langle I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\succ 0$ for all ${\boldsymbol{x}}\in\overline{\mathcal{X}}$. Since $\overline{\mathcal{X}}$ is compact and $F({\boldsymbol{x}})\succ 0$ on ${\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, letting
$\varepsilon\coloneqq\frac{\min_{{\boldsymbol{x}}\in\overline{\mathcal{X}}}\lambda_{\min}(F({\boldsymbol{x}}))}{2\max_{{\boldsymbol{x}}\in\overline{\mathcal{X}}}|\lambda_{\max}\left(I_{p}+\int_{\mathcal{Y}}\left\langle I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)|},$ (24)
we obtain the desired conclusion. ∎
We now prove Theorem 4.2.
Proof of Theorem 4.2.
Let $F_{\varepsilon}({\boldsymbol{x}})$ be the polynomial matrix in Lemma 6.2 where $\varepsilon>0$ is defined in (24). Then, applying Theorem 6.1 to $F_{\varepsilon}({\boldsymbol{x}})$ yields $N_{1}\in\mathbb{N}$ and polynomial matrices $^{\prime}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$^{\prime}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in $\boldsymbol{x}$} and satisfy that $S^{\prime}_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$, $S^{\prime}({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$’s are SOS matrices, such that
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{1}}F_{\varepsilon}({\boldsymbol{x}})=S^{\prime}_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S^{\prime}({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Then, by the definition of $F_{\varepsilon}({\boldsymbol{x}})$,
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg F+N_{1}}F({\boldsymbol{x}})=\Gamma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Gamma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),$ (25)
where
$\Gamma_{0}({\boldsymbol{x}})\coloneqq\varepsilon({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D+N_{1}}I_{p}+S^{\prime}_{0}({\boldsymbol{x}}),\ \ \Gamma({\boldsymbol{x}},{\boldsymbol{y}})\coloneqq\varepsilon({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg_{{\boldsymbol{x}}}G+N_{1}}I_{pq}+S^{\prime}({\boldsymbol{x}},{\boldsymbol{y}}).$
Note that $\Gamma_{0}({\boldsymbol{x}})$ and $\Gamma({\boldsymbol{x}},{\boldsymbol{y}})$ are homogeneous in $\boldsymbol{x}$} because by Theorem 6.1,
$\deg S^{\prime}_{0}=N_{1}+\deg F_{\varepsilon}=D+N_{1}\ \ \text{and}\ \ \deg_{{\boldsymbol{x}}}S^{\prime}=N_{1}+\deg F_{\varepsilon}-\deg_{{\boldsymbol{x}}}G=D-\deg_{{\boldsymbol{x}}}G+N_{1}.$
Since $S^{\prime}_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$, $S^{\prime}({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$ are SOS matrices, for all ${\boldsymbol{x}}\in\{{\boldsymbol{x}}\in{\mathbb{R}}_{+}^{n}\mid{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}}=1\}$, it holds
$\Gamma_{0}({\boldsymbol{x}})\succeq\varepsilon I_{p}\ \text{ and }\ \Gamma({\boldsymbol{x}},{\boldsymbol{y}})\succeq\varepsilon I_{pq},\ \forall{\boldsymbol{y}}\in\mathcal{Y}.$
Write $\Gamma({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\alpha}}\in\mbox{supp}_{{\boldsymbol{x}}}(\Gamma)}\Gamma_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ and define
$L(\Gamma)\coloneqq\max_{{\boldsymbol{y}}\in\mathcal{Y}}\max_{{\boldsymbol{\alpha}}\in\mbox{supp}_{{\boldsymbol{x}}}(\Gamma)}\frac{{\boldsymbol{\alpha}}!}{\deg_{{\boldsymbol{x}}}\Gamma!}\|\Gamma_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\|.$
As $\mathcal{Y}$ is compact, $L(\Gamma)$ is well-defined. Let
$N_{2}\coloneqq\max\left\{\frac{\deg\Gamma_{0}(\deg\Gamma_{0}-1)L(\Gamma_{0})}{2\varepsilon}-\deg\Gamma_{0},\ \frac{\deg_{{\boldsymbol{x}}}\Gamma(\deg_{{\boldsymbol{x}}}\Gamma-1)L(\Gamma)}{2\varepsilon}-\deg_{{\boldsymbol{x}}}\Gamma\right\}.$
Let $P_{0}({\boldsymbol{x}})=({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{2}}\Gamma_{0}$ and $P({\boldsymbol{x}},{\boldsymbol{y}})=({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{2}}\Gamma$. Write
$P_{0}({\boldsymbol{x}})=\sum_{|{\boldsymbol{\alpha}}|=\deg P_{0}}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\quad\text{ and }\quad P({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}P}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}.$
By Theorem 6.2, $P_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg P_{0}$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}P$ and ${\boldsymbol{y}}\in\mathcal{Y}$. Letting $N=D-\deg F+N_{1}+N_{2}$, the conclusion follows from (25). ∎
7. Conclusions
We extend various classical Positivstellensätze to provide SOS-structured characterizations for polynomial matrices that are positive (semi)definite over a semialgebraic set defined by a PMI with UQs. Under the Archimedean condition, we first present a matrix-valued Positivstellensatz incorporating universal quantifiers, along with a sparse version for scalar-valued polynomial objectives, leveraging the correlative sparsity patterns. We also derive two generalized Positivstellensätze without assuming the Archimedean condition. These results significantly extend the existing work on Positivstellensätze, and some of them remain novel and intriguing even in the absence of UQs. Our results are valuable in ensuring the robustness of PMIs over a prescribed set with uncertainty, allowing potential applications in areas such as optimal control, systems theory, and certifying stability.
Acknowledgements
The authors appreciate Professor Konrad Schmüdgen for providing the result in [44, Theorem 10.25], which significantly simplified the proof of Theorem 3.1. Feng Guo was supported by the Chinese National Natural Science Foundation under grant 12471478. Jie Wang was supported by National Key R&D Program of China under grant No. 2023YFA1009401, the Strategic Priority Research Program of the Chinese Academy of Sciences XDB0640000 & XDB0640200, and the National Natural Science Foundation of China under grant No. 12201618 & 12171324.